### v2

# Taxonomy of differentiationP. Komarov, F. van Breugel, J. N. Kutz

Abstract
Differentiation is a cornerstone of computing and data analysis in every discipline of science and engineering. Indeed, most fundamental physics laws are expressed as relationships between derivatives in space and time. However, derivatives are rarely directly measurable and must instead be computed, often from noisy, potentially corrupt data streams. There is a rich and broad literature of computational differentiation algorithms, but many impose extra constraints to work correctly, e.g. periodic boundary conditions, or are compromised in the presence of noise and corruption. It can therefore be challenging to select the method best-suited to any particular problem. Here we review a broad range of numerical methods for calculating derivatives, present important contextual considerations and choice points, compare relative advantages, and provide basic theory for each algorithm in order to assist users with the mathematical underpinnings. This serves as a practical guide to help scientists and engineers match methods to application domains. We also provide an open-source Python package, PyNumDiff, which contains a broad suite of methods for differentiating noisy data.
keywords:
numerical differentiation, derivatives, time series data, data science
{AMS}
97M01 (Math Ed: Mathematical modeling, applications of mathematics: Instructional exposition / tutorial papers), 37M01 (Dynamical Systems: Approximation methods and numerical treatment of dynamical systems: Instructional exposition / tutorial papers)
Contents
1 Introduction
Computing derivatives is a critically enabling task for almost any application in the engineering, physical, mathematical, and biological sciences. Mathematical models and governing equations often take the form of relationships between changing quantities, so whether the computation is of temporal or spatial data, simulated or measured, be it in the field of signal processing, system identification, or Machine Learning (ML), accurate derivative estimates are crucial. The great need in this area has fueled great innovation, resulting in a proliferation of differentiation methods specialized to take advantage of certain scenarios or system properties [savitzky1964smoothing, trefethen2000spectral, kutzbook, aravkin]. However, confronting so many options complicates algorithm selection.
Numerical algorithms including Finite Differences, Spectral Methods, and Finite Elements or Volumes are now mature technologies, with the latter powering commercialized (e.g. COMSOL [comsol], Ansys and Abaqus) and open-source (e.g. FEniCS, Firedrake, OpenFOAM) PDE solvers. Likewise, Automatic Differentiation (AutoDiff) dominates in Machine Learning [jax2018github]. As practitioners in the rapidly-growing field of data-driven science, we also take particular interest in derivative estimation for noisy data streams. Motivated by the definition of the derivative from calculus, beginners in numerical computation often use finite differences by default, but this turns out to be one of the worst choices for real data. We attempt to give comprehensive coverage to all these schemes, from the narrowly focused to the most general-purpose. The overall trend is that more constrictive situations allow for better performance, so we suggest the use of specialized tools if possible. Or, if faced with a dearth of simplifying assumptions, we have made it easy for the reader to skip ahead to the pertinent category.
This paper is organized as follows: In section 2 we establish a common map for navigating key determining factors, so practitioners can quickly get a sense of which archetypal situation is most representative of their differentiation problem and see at a glance which solution strategies might work best. Then we follow that map, beginning with the easiest situations and most bulletproof methods, working toward more equivocal, ambiguous scenarios. In section 3 we cover Automatic Differentiation tools like JAX [jax2018github], which have been incredibly successful and revolutionized deep learning, but are by nature only narrowly applicable. We also touch on their use in Differentiable Physics simulators. In section 4 we address noiseless simulation, including Finite Differencing, Spectral Methods, and Finite Elements. We then discuss noise as a concept in section 5 and follow up in section 6 with methods that can address uncertainty by taking advantage of additional system knowledge in the form of dynamics models and synchronized data streams. The more difficult case of differentiating in the presence of noise without additional information is covered in section 7, where we describe an approach for measuring performance in the absence of ground truth by balancing fidelity and smoothness [floris2020], use this metric to choose near-optimal hyperparameters, and present experimental results to compare accuracy and bias across methods. In doing so, we make heavy use of tools from our companion, open-source Python package PyNumDiff [pynumdiff], which implements many of the methods discussed as well as hyperparameter optimization to tune them. In section 8 we cover the practical consideration of data with irregular sampling rate. Finally, section 9 concludes with general retrospective and recommendations.
2 Choosing a Method
[…]
Figure 1: Preferred differentiation algorithms at a glance, for the five major situations identified in this review. Names are shorthand, with details given in later sections, hyperlinked for convenience. Further details and rationale behind these distinctions and selections are mapped in Figure 2. $\Delta x$ and $\Delta t$ are spacing between samples, $m$ is FD scheme order, $h$ is element side-length, and $\alpha$ is a constant, making both the FEM and FD error bounds similarly algebraic, while error of Spectral Methods decreases “super-algebraically” as the number of samples and basis functions increases.
Start static, analytic? Automatic Differentiation noiseless simulation? smooth derivative? regular domain? smooth derivative? periodic BC? Can’t distinguish noise from derivative Finite Elements Fourier
Spectral
can sample cos-spaced? Chebyshev
Spectral
Finite Difference periodic BC? Fourier Spectral known dynamics? linear? Gaussian white noise? Kalman Smoother Robust Kalman Smoother Nonlinear Kalman Smoother irregular steps? Kernel Smoothing
$\rightarrow$ Finite Difference
Iterated Finite Difference Polynomial Fits Basis Fits
with Tricks
Total Variation Regularized Derivative Kalman Smoothing with Naive Model
yes no, dynamic or non-analytic solutions yes, governing equations no, measurementsnonoyesyesnoyes noisy signal,
smooth derivative
yesnoyesno, uniform samplesyesnoyesyesyesno classic often works nono modeling may not be worth it noyes
Figure 2: Suggested procedure for choosing a differentiation method. Methods are boxed with blue, violet, purple, or red to give an overall impression of how computationally efficient, accurate, robust, and user-friendly they are, with blue meaning a method excels in all respects for its specialization and red meaning a solution is hopeless. Groupings in the lowermost node represent many methods, expounded in section 7. Methods and a couple of other labels are hyperlinked for quick navigation to relevant sections.
In the broadest possible strokes, there are three scenarios for computing derivatives, two of which can be further refined:
  1. 1.
    Analytic relationships with unchanging structure
  2. 2.
    Simulations: (a) less complicated, (b) more complicated
  3. 3.
    Noisy data: (a) with a model, (b) without a model
A digest of these with some preferred algorithms is given in Figure 1. Figure 2 shows a more complete and detailed overview of key considerations and methods, organized as a flowchart. The decisions therein are important because different techniques make different assumptions. For example, smoothing methods assume the signal is continuously differentiable up to some order (possibly $C^{\infty}$), and the Kalman filter requires a system model. Under changing assumptions, the most suitable approach changes, so it is important to understand whether data is periodic, noisy, etc. Notably missing among these considerations is dimensionality: Because differentiation is a linear operator, all methods extend equally well to the multidimensional case, either by multivariable chain rule, by using higher-dimensional meshes and basis functions, or by sequential application along data dimensions (in any order). We therefore primarily treat the one-dimensional case in this review, without loss of generality.
For most differentiation scenarios (sections 3, 4, 6), there are only a handful of clear winners that dominate other methods due to computational efficiency, accuracy, or the unique ability to handle relevant structure. The exceptional case is differentiation of noisy data without a model (section 7), because there are many possible underlying generating functions, making the problem ill-posed. Choosing a method becomes significantly more complicated in this situation, because so many methods have been invented, none foolproof. Describing the mathematical underpinnings of all these methods, as well as the problems they fit, requires an extensive vocabulary of symbols, so a partial lexicon is given in Table 1 to aid with orientation.
Table 1: Common symbols used throughout this review.
$y$, $u$, $v$ a function, although $y$ can also be a spatial independent variable
$f$ a function, also used to denote frequency
$x$, $t$, $\Delta x$, $\Delta t$ independent variable and its discrete increment
$c$, $a$, $b$, $\alpha$, $\beta$ constants or coefficients
$\epsilon$, $h$ small values
$n$ data sample iterator
$k$ basis function or term iterator
$j$ generic iterator
NN number of total data samples
$\xi$, $\varphi$ basis functions
[…]
$\omega$ angular frequency
$\boldsymbol{\theta}$ parameters, stacked together as a vector
$X$ input data examples stacked together in a matrix or tensor
$Y$ target examples stacked together, or DFT coef. with subscript $k$
$L$ a loss function, or occasionally the length of an interval
2.1 Recommendations
When suitable, Automatic Differentiation (section 3) and Spectral Differentiation with the Fourier (section 4.2.1) or Chebyshev basis (section 4.2.2) are absolute best-in-class, but use of these is quite restrictive. The Finite Element Method (footnote 5), by contrast, is exceptionally versatile, capable of differentiating even non-smooth functions over irregularly-shaped domains. But it is also rather mathematically involved, involves more setup, and is relatively computationally demanding—although much of its administrative overhead can be offloaded to commercial or open-source packages. Finite Difference (section 4.1) is less accurate and robust but is very straightforward, works well in the absence of noise, and has well-defined error bounds [kutzbook].
Down the “noisy signal, smooth derivative” path, the Fourier spectral method (i.e. transform, zero out higher modes, and inverse transform) is ideal for isolating high-frequency noise (section 5), but it should only be used when signals are periodic (can be joined smoothly end-to-end), lest Gibbs phenomenon (Figure 7) corrupt the approximation. Otherwise, one should exploit a model of system dynamics and noise when possible (section 6), although system models can be challenging to derive beyond simple systems and can be of limited utility unless fairly accurate. In the naive case (section 7), sophisticated methods tend to perform similarly (section 7.8), but additional considerations—such as robustness to outliers, fast hyperparameter optimization, and ability to handle large or variable step size—distinguish a few methods for specialized use (Figure 57). In some cases, a method may even be preferred for narrative reasons, having theoretical structure that matches assumptions about an underlying process or context.
3 Differentiating Analytic Functions with Static Structure
Given the dramatic success of deep learning, optimized by hardware-accelerated gradient descent implemented with libraries like JAX [jax2018github] and PyTorch [pytorch], it is natural to wonder how AutoDiff fits into the larger numerical differentiation ecosystem. In a word: AutoDiff is not for differentiating data samples, but rather for sampling fixed derivative relationships. Automatic Differentiation is most commonly used in the context of machine learning, where a model, typically named $f$ for “function”, is parameterized by a big vector $\boldsymbol{\theta}$, takes a batch of inputs $X$, and produces a batch of outputs.
The paradigmatic supervised ML problem is [^1] A computer scientist we know says, “Supervised learning is the only thing that really works.” [abhishek] There are of course other kinds of problems, such as unsupervised learning, semi-supervised learning, self-supervised learning [barlowtwins], and reinforcement learning, but ultimately all are reduced to minimizing loss based on known correct answers or clever assumptions that enforce self-consistency and regularization [sindy-shred, parsimony].:
$\boldsymbol{\theta}^{*}=\operatorname*{argmin}_{\theta\in\mathbb{R}^{m}}\ L(f(X;\boldsymbol{\theta}),Y)$
where $L$ is a loss function, minimized when model outputs equal targets; $f$ is a model with possibly multidimensional input and output; $X$ is a set of inputs to the model; $Y$ is a set of corresponding target outputs; and $\boldsymbol{\theta}$ affects how the model produces its answers, sometimes shown after a semicolon for visual separation from model inputs. The key realization of ML is that we can consider $X$ and $Y$ to be static and given, and treat the model’s parameterization, $\boldsymbol{\theta}$, as an input, thus obtaining a relationship between overall error and the parameters. If we then differentiate this relationship with respect to the parameters and evaluate at the given input data, we obtain a gradient leading toward the loss function’s minimum, which we can use to update model parameters as:
$\boldsymbol{\theta}_{j+1}=\alpha\boldsymbol{\theta}_{j}-\epsilon\nabla_{\boldsymbol{\theta}}L\Big|_{X,Y;\boldsymbol{\theta}_{j}}$
where $\alpha$ is a momentum parameter; $\epsilon$ is a learning rate; and $\nabla_{\boldsymbol{\theta}}L\big|_{X,Y;\boldsymbol{\theta}_{j}}$ indicates the gradient of $L$, using the current parameterization of $f$, averaged over evaluations with input-output data pairs [goodfellow].
The total size of the data may be massive, so we often choose to evaluate over small sub-samples of the data for computational efficiency, a practice known as stochastic gradient descent. This leads to “sampling noise”, because the mean gradient over the sub-population is not exactly the same as the mean gradient evaluated over the full population. To handle this noise, we may use a more sophisticated update rule, like Adam [adam-optimizer], which effectively serves as a low-pass filter to smooth out variations in the descent direction. For more on noise, see section 5.
To get a better sense of what taking the gradient of such a complicated construction means, we can expand the calculation using the multivariable chain rule:
$\nabla_{\boldsymbol{\theta}}L=\frac{d}{d\boldsymbol{\theta}}L(f(\mathbf{x};\boldsymbol{\theta}),\mathbf{y})=\frac{\partial L}{\partial f}\frac{df(\mathbf{x};\boldsymbol{\theta})}{d\boldsymbol{\theta}}+\overset{\text{\normalsize 0}}{\cancel{\frac{\partial L}{\partial\mathbf{y}}\frac{d\mathbf{y}}{d\boldsymbol{\theta}}}}=\frac{\partial L}{\partial f}\Big(\overset{\text{\normalsize 0}}{\cancel{\frac{\partial f}{\partial\mathbf{x}}\frac{d\mathbf{x}}{d\boldsymbol{\theta}}}}+\frac{\partial f}{\partial\boldsymbol{\theta}}\overset{\text{\normalsize 1}}{\cancel{\frac{d\boldsymbol{\theta}}{d\boldsymbol{\theta}}}}\Big)=\frac{\partial L}{\partial f}\frac{\partial f}{\partial\boldsymbol{\theta}}$
If the model’s evaluation function is compositional, as in the layers of a deep net, we perform the chain rule through all the layers. This involves differentiating vectors, matrices, or even tensors w.r.t. other multidimensional objects, which can be somewhat confusing but is ultimately a pile of scalar derivatives stacked together. The details are well explained by [erik] and [clark]. The final product of this process is a relationship in terms of the loss function and model evaluation function, both of which are fixed. AutoDiff exploits these unchanging relationships to find the analytic derivative (in many variables) of an analytic function in a very mechanistic way (computational graph, Figure 3), then evaluates it at many data points in parallel to arbitrarily high accuracy.
[…]
Figure 3: Example of a simple computational graph for $\nabla_{\boldsymbol{\theta}}L$ with compositional $f(\mathbf{x};\boldsymbol{\theta})=f_{3}(f_{2}(f_{1}(\mathbf{x};\boldsymbol{\theta}_{1});\boldsymbol{\theta}_{2});\boldsymbol{\theta}_{3})$, which results in $\frac{\partial f}{\partial\boldsymbol{\theta}}=\frac{\partial f_{3}}{\partial f_{2}}\frac{\partial f_{2}}{\partial f_{1}}\frac{\partial f_{1}}{\partial\boldsymbol{\theta}_{1}}+\frac{\partial f_{3}}{\partial f_{2}}\frac{\partial f_{2}}{\partial\boldsymbol{\theta}_{2}}+\frac{\partial f_{3}}{\partial\boldsymbol{\theta}_{3}}$ by chain rule.
3.1 When Automatic Differentiation is Inappropriate
To get a sense of why such a dominant tool as AutoDiff does not comprehensively handle our differentiation needs, consider simulating the 2D wave equation [trefethen2000spectral], which obeys the partial differential equation (PDE):
$\frac{\partial^{2}u}{\partial t^{2}}=\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}$
where $x$ and $y$ are dimensions of the input, and $u(x,y)$ is the wave amplitude. We choose the domain $1\leq x,y\leq 1$, boundary condition $u=0$, and initial condition $u(x,y,0)=e^{-40((x-0.4)^{2}+y^{2})},\ \frac{\partial}{\partial t}u(x,y,0)=0$, which evolves as illustrated in Figure 4 through time.
Figure 4: Simulation of the wave equation for initial condition $u(x,y,0)=e^{-40((x-0.4)^{2}+y^{2})}$, $u_{t}(x,y,0)=0$ and boundary condition $u=0$.
To efficiently simulate the evolution of this system, we can discretize and use 1, which employs the “leapfrog” scheme, known to be stable for the second-order wave equation with small enough $\Delta t$ [strang-notes, kutzbook]. At each time step, $u$ is updated by accumulating small numerical integrals of its second-order spatial derivatives. This means that although the initial condition of $u$ is simply analytic, simulated $u$ rapidly acquires more terms across future time steps. Furthermore, for arbitrary initial and boundary conditions, it is possible an analytic expression for $u$ does not exist.
Algorithm 1 Wave Simulation
1:choose final time $T$, discrete time increment $\Delta t$, and spatial gaps $\Delta x$ and $\Delta y$
2:$u=$ sample the initial condition on a grid, where samples are $\Delta x$ and $\Delta y$ apart in respective dimensions
[…]
4:for $t_{n}\in\{0,\Delta t,2\Delta t,...,T\}$ do
5:  $\text{right-hand side}=\Big[\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}\Big]\Big|_{\text{grid points}}$ $\triangleright$ evaluate at grid points
6:  $u_{\text{next}}=2u-u_{\text{prev}}+\Delta t^{2}\cdot\text{(right-hand side)}$ $\triangleright$ leapfrog update
7:  $u_{\text{next}}[\text{boundary}]=0\triangleright$ pin the boundary to 0
8:  $u_{\text{prev}}=u$
9:  $u=u_{\text{next}}$
10:end for
Notice that the derivatives here are taken on a changing function at fixed sample points rather than on a fixed relationship at changing parameterization points. This distinction is subtle, because a parameterized relationship with sufficient expressive power can model a function of any shape [hornik1989], so in principle we could define $u(x,y,t;\boldsymbol{\theta})$ to follow a changing function and use AutoDiff to evaluate its derivatives w.r.t. $x$ and $y$ at space-time grid points. But this would require knowledge of the true shape of $u(x_{n},y_{n})\ \forall\ t_{n}$ to use as a training target, and the whole point of simulation is to compute these answers using only the PDE relationship, without having to solve analytically or encode $u$ as a functional relationship when numerical values will do.
3.2 Automatically Differentiable Physics
There is a case where AutoDiff can play an important role in simulations: If we assume a particular, unchanging form between a function’s values and those of its derivative, like a Finite Difference equation (see section 4.1) or FFT-based Spectral Method (see 2 and 3), then Automatic Differentiation can be used to take gradients of this relationship, a kind of meta-derivative.
Several software libraries implement this architecture, including JAX-Fluids [jaxfluids-article, jaxfluids-library], PhiFlow [phiflow], JAX-FEM [jax-fem], and Exponax [exponax]. “Differentiable Physics” [differentiable-physics] refers to the passage of gradients all the way through a physics simulation, enabling seamless assimilation in ML training and inference pipelines. Note, however, that the physical derivatives per se of such simulations are still found via whichever proximal differentiation algorithm the simulator implements (Finite Difference in the case of [jaxfluids-library, phiflow]), and so the simulation itself is only as accurate as that method rather than inheriting AutoDiff’s fidelity (down to numerical precision).
Figure 5: From [differentiable-physics], used with permission, a neural network’s solution to an advection flow inverse problem, demonstrating the strength of end-to-end training with Differentiable Physics. Fluid with marker dye is simulated at a particular location, and the goal is to control high-dimensional fluid flows to force the marker patch into a target shape.
3.3 Recommendations
In some cases it may be possible to cast a problem into a static, analytic form. If so, use Automatic Differentiation without reservation, knowing it has been stress tested and scaled up by the prolific Machine Learning community. But recognize these are constrictive limitations, that for many problems analytic forms are unknown or inconvenient, necessitating a purely numeric representation of functions and other differentiation methods.
4 Differentiating Noiseless Simulation Data
When a function $y(x)$ ceases to be representable with a few analytic terms, it is best represented numerically, by its values. Because continuous function values are infinitely dense, the only way to practically do this is with discrete samples, $y_{n}=y(x_{n}),\ n\in\{0,...,N-1\}$, for a total of $N$ points over some domain, $x\in[a,b]$. In this situation, taking derivatives becomes less exact because isolated samples do not allow evaluation of slope in the limit as $dx\to 0$. Instead, we remember samples are taken from some continuous underlying function, assume a particular interpolation to fill the gaps between samples, and differentiate that interpolation [johnson-fft-notes]. Fortunately, in the absence of noise, these approximations can be made highly accurate.
4.1 Finite Difference
Consider the most basic definition of a derivative, which takes the difference of two nearby function values, divides by the distance between them, and shrinks that distance to an infinitesimal value:
$y^{\prime}(x)=\lim_{h\to 0}\frac{y(x+h)-y(x)}{h};\quad y^{\prime}_{n}=\lim_{\Delta x\to 0}\frac{y_{n+1}-y_{n}}{\Delta x}$
On the left is the classical continuous definition, based on a linear interpolation, and on the right is the discrete analog with $y_{n}=y(x_{n})$, which is the discrete version. Clearly, if we could sample a function infinitely densely, this would evaluate the derivative correctly.
But in computation, the limit can never be perfectly evaluated numerically; rather $\Delta x$ is taken to be sufficiently small so the approximation is accurate to within some acceptable error. For smooth functions, we can bound this error by substituting a truncated Taylor series, which effectively interpolates the function by substituting a degree-$d$ polynomial (in $\Delta x$) for $y_{n+1}$, e.g. for $d=2$:
$\frac{y_{n+1}-y_{n}}{\Delta x}=\frac{y_{n}+\Delta x\cdot y^{\prime}_{n}+\frac{\Delta x^{2}}{2!}y^{\prime\prime}(c)-y_{n}}{\Delta x}=y^{\prime}_{n}+\frac{\Delta x}{2!}y^{\prime\prime}(c)$
for some $c\in[x,x+\Delta x]$. Because $\frac{y^{\prime\prime}(c)}{2!}$ is simply a constant, we can say this differencing equation is within $O(\Delta x)$ of the true sampled derivative. [^2] The use of big-O notation is common with Taylor series, but it is technically only allowed in cases where the remainder term can be given an explicit form based on Intermediate Value Theorem. If the function under consideration, $y$, is not continuously differentiable $d$ times on the closed interval $x,x+\Delta x]$, then the error of the numerator Taylor expansion (before division by $\Delta x$) is $o(\Delta x^{d-1})$ rather than $O(\Delta x^{d})$ [little-o-big-o]. Little-o, “$f$ is o(g)o(g)”, means $\lim_{N\to\infty}\frac{f(N)}{g(N)}=\lim_{\Delta x\to 0}\frac{f(\Delta x)}{g(\Delta x)}=0$, while big-O, “$f$ is O(g)O(g)”, means $|f(N)|\leq c|g(N)|\ \forall\ N\geq N_{0}\Leftrightarrow|f(\Delta x)|\leq c|g(\Delta x)|\ \forall\ \Delta x\leq\Delta x_{0}$ for some $c>0$ and $N_{0}$ or $\Delta x_{0}\in\mathbb{R}$. For $g$ of the same degree, little-o is (perhaps counterintuitively) a stronger statement than big-O, but little-o of one degree down is a weaker statement. At this place in the flowchart (Figure 2), we are considering smooth derivatives, so we assume infinite continuous differentiability, $y\in C^{\infty}$, and adopt big-O. This bound can be shrunk by using more points from a “stencil”, like [1,0,1][-1,0,1] (i.e. the previous point, current point, and following point, “centered” on the current point), and setting up a system of equations to cleverly cancel terms:
(1) $y^{\prime}_{n}\approx$ $\ c_{0}y_{n-1}+c_{1}y_{n}+c_{2}y_{n+1}\quad\quad\text{(ansatz)}$
$=$ $\ c_{0}[y_{n}-\Delta x\cdot y^{\prime}_{n}+\frac{(-\Delta x)^{2}}{2!}y^{\prime\prime}_{n}+O(\Delta x^{3})]+c_{1}y_{n}$
[…]
By grouping terms according to derivative order, we can obtain as many equations as unknown coefficients:
$y_{n}:\quad 0$ $=c_{0}+c_{1}+c_{2}$
$y^{\prime}_{n}:\quad 1$ $=-\Delta x\cdot c_{0}+\Delta x\cdot c_{2}$
$y^{\prime\prime}_{n}:\quad 0$ $=\frac{(-\Delta x)^{2}}{2!}c_{0}+\frac{\Delta x^{2}}{2!}c_{2}\rightarrow 0=c_{0}+c_{2}$
These can be set up as a linear inverse problem:
$\begin{bmatrix}1&1&1\\ -1&0&1\\ 1&0&1\end{bmatrix}\begin{bmatrix}c_{0}\\ c_{1}\\ c_{2}\end{bmatrix}=\begin{bmatrix}0\\ \frac{1}{\Delta x}\\ 0\end{bmatrix}$
and solved to yield $c_{0}=\frac{-1}{\Delta x},\ c_{1}=0,\ c_{2}=\frac{1}{\Delta x}$. Plugging these back into Equation 1, the $\Delta x$ in the denominators cancels one of those in the error terms, making the final error $O(\Delta x^{2})$.
This methodology can be generalized [taylor-fd] to higher, $\nu^{\text{th}}$-order, derivatives and arbitrary stencils of length $S$, $s_{0},...,s_{S-1}]$, including lopsided, one-sided (“forward” or “backward” difference), and non-contiguous ones; the linear inverse problem just becomes:
(2) $\begin{bmatrix}s_{0}^{0}&\cdots&s_{S-1}^{0}\\ \vdots&\ddots&\vdots\\ s_{0}^{S-1}&\cdots&s_{S-1}^{S-1}\end{bmatrix}\begin{bmatrix}c_{0}\\ \vdots\\ c_{S-1}\end{bmatrix}=\begin{bmatrix}0\\ \vdots\\ \frac{\nu!}{\Delta x^{\nu}}\\ \vdots\\ 0\end{bmatrix}\leftarrow\parbox{56.9055pt}{nonzero only\\ at $\nu^{\text{th}}$ index}$
The error of the resulting finite difference scheme, $y^{\nu}_{n}\approx c_{0}y_{n+s_{0}}+...+c_{S-1}y_{n+s_{S-1}}$, is at most $O(\Delta x^{S-\nu})$, because plugging coefficients back into the scheme’s Taylor-expansion causes all terms up to and including the $(S-1)^{\text{th}}$ order—except the $\nu^{\text{th}}$—to cancel by construction, leaving only $O(\Delta x^{S})$ beyond, which is then divided by $\Delta x^{\nu}$}. In the special case of even $\nu$ and an odd-length, centered stencil, the solution causes cancellation of all odd-order terms, including the most dominant error term, $O(\Delta x^{S})$, allowing us to achieve higher accuracy with the same number of points. Table 2 gives formulas with second-order accuracy. For greater accuracy, one can solve Equation 2 with longer stencils, i.e. more neighboring points.
Table 2: Second-order-accurate center-difference formulas along with forward- and backward-difference schemes for endpoints.
$\mathbf{O(\Delta x^{2})}$ center-difference schemes
$y^{\prime}_{n}=[y_{n+1}-y_{n-1}]/2\Delta x$
$y^{\prime\prime}_{n}=[y_{n+1}-2y_{n}+y_{n-1}]/\Delta x^{2}$
$y^{\prime\prime\prime}_{n}=[y_{n+2}-2y_{n+1}+2y_{n-1}-y_{n-2}]/2\Delta x^{3}$
$y^{(4)}_{n}=[y_{n+2}-4y_{n+1}+6y_{n}-4y_{n-1}+y_{n-2}]/\Delta x^{4}$
$\mathbf{O(\Delta x^{2})}$ forward- and backward-difference schemes
$y^{\prime}_{n}=[-3y_{n}+4y_{n+1}-y_{n+2}]/2\Delta x$
$y^{\prime}_{n}=[3y_{n}-4y_{n-1}+y_{n-2}]/2\Delta x$
$y^{\prime\prime}_{n}=[2y_{n}-5y_{n+1}+4y_{n+2}-y_{n+3}]/\Delta x^{2}$
$y^{\prime\prime}_{n}=[2y_{n}-5y_{n-1}+4y_{n-2}-y_{n-3}]/\Delta x^{2}$
Finite difference formulas can be compactly represented as sparse matrices acting on a vector of data by simple multiplication. For example, the following two matrices compute the first and second derivatives of data vector $\mathbf{y}$ using the formulas from Table 2. Center differencing is used for interior points, but the first and last rows reflect forward and backward difference formulas, respectively.
$\mathbf{y}^{\prime}=\frac{1}{2\Delta x}\begin{bmatrix}-3&4&-1&0&\cdots&0\\ -1&0&1&0&\cdots&0\\ 0&-1&0&1&\cdots&0\\ \ddots&\ddots&\ddots&\ddots&\ddots&\ddots\\ 0&\cdots&0&-1&0&1\\ 0&\cdots&0&1&-4&3\end{bmatrix}\begin{bmatrix}y_{1}\\ y_{2}\\ \vdots\\ y_{n}\end{bmatrix}={\bf D}\cdot\mathbf{y}$
$\mathbf{y}^{\prime\prime}=\frac{1}{\Delta x^{2}}\begin{bmatrix}2&-5&4&-1&0&\cdots\\ 1&-2&1&0&0&\cdots\\ 0&1&-2&1&0&\cdots\\ \ddots&\ddots&\ddots&\ddots&\ddots&\ddots\\ \cdots&0&0&1&-2&1\\ \cdots&0&-1&4&-5&2\end{bmatrix}\begin{bmatrix}y_{1}\\ y_{2}\\ \vdots\\ y_{n}\end{bmatrix}={\bf D_{2}}\cdot\mathbf{y}$
Finite Differences are local approximations, using only a small set of neighboring data points to estimate derivatives. The method is typically constructed to use the minimal number of neighbors to achieve $O(\Delta x^{m})=O(N^{-m})$ error, where $m$ is called the “order” [kutzbook] and $N$ is the number of samples, because $\Delta x\propto 1/N$ on a fixed domain.
4.1.1 Recommendations
Finite Difference is simple and flexible. It realizes a relationship between function and derivative values that is at once easy to implement and understand, which has made it the method of choice for several meta-differentiation libraries (section 3.2). It works with arbitrary boundary conditions and is fast to run, $O(Nm)$, requiring only one major iteration along the data. However, Finite Difference’s error bounds only hold for smooth functions and in the absence of noise, the same situation where Spectral Methods, which achieve better accuracy, can excel, so Finite Difference should only be used if Fourier (section 4.2.1) or Chebyshev (section 4.2.2) do not apply.
4.2 Spectral Methods
This approach forms a continuous interpolant, $y(x)$, from a set of “basis functions”, $\{\xi_{k}(x)\},\ k\in\mathbb{N}$, defined over the domain, $x\in[a,b]$. This $y$ can then be differentiated and sampled. The basis is chosen to span a function space containing the function of interest, meaning $y$ can be represented as a linear combination of the basis functions:
(3) $y(x)=\sum_{k}c_{k}\xi_{k}(x),\quad x\in[a,b]$
where each $c_{k}$ is a scalar coefficient. A basis set (e.g. the Fourier functions in the next section) can have infinitely many members, and we may need them all to fully represent an arbitrary function. But in computation, we always truncate to consider only a finite number of terms, often $N=|\{y_{n}\}|$, so $y$ has just enough expressive power to uniquely interpolate $N$ samples.
Members of a basis set are linearly independent, i.e. not mere multiples of one another, but often basis functions are chosen to meet the stronger condition of orthogonality, meaning the inner product between them is defined:
$\langle\xi_{i},\xi_{j}\rangle=\int\limits_{a}^{b}\xi_{j}(x)\overline{\xi_{i}(x)}dx=\begin{cases}0&i\neq j\\ \text{constant}\in\mathbb{R}&i=j\end{cases}$
where the overbar denotes a complex conjugate (for complex-valued basis functions).
Taking the inner product of both sides of Equation 3 against the $k^{\text{th}}$ orthogonal basis function, all terms of the sum but the $k^{\text{th}}$ disappear, and we can rearrange to find the coefficient, $c_{k}$:
(4) $\frac{\langle\xi_{k},y\rangle}{\langle\xi_{k},\xi_{k}\rangle}=\frac{\int\limits_{a}^{b}y(x)\overline{\xi_{k}(x)}dx}{\|\xi_{k}(x)\|_{2}^{2}}=c_{k}$
Note that none of the inner product integrals here are allowed to diverge, which limits $\{\xi_{k}\}$ and $y$ to the function space $L_{2}$, the square integrable functions, by definition. This is the only Lebesgue space which is also a Hilbert space (stronger condition), due to the presence of the inner product. This means functions in $L_{2}$ can be treated completely analogously to vectors, so finding $\{c_{k}\}$ can be thought of as similar to a vector change of basis, as shown in Figure 6.
$\vec{e}_{0}$$\vec{e}_{1}$$\vec{y}$$\vec{\xi}_{0}$$\vec{\xi}_{1}$$\mathrm{proj}_{\vec{\xi}_{0}}\vec{y}$$\mathrm{proj}_{\vec{\xi}_{1}}\vec{y}$
Figure 6: A vector change of coordinates, expressing vector $\vec{y}$ in terms of the spanning basis vectors $(\vec{\xi}_{0},\vec{\xi}_{1})$ instead of axis-aligned unit vectors $(\vec{e}_{0},\vec{e}_{1})$, can be accomplished by $\vec{y}=\alpha\vec{\xi}_{0}+\beta\vec{\xi}_{1}$. If $\vec{\xi}_{0}$ and $\vec{\xi}_{1}$ are orthogonal, then $\alpha=\frac{\langle\vec{\xi_{0}},\vec{y}\rangle}{\|\vec{\xi}_{0}\|_{2}^{2}},\ \beta=\frac{\langle\vec{\xi}_{1},\vec{y}\rangle}{\|\vec{\xi}_{1}\|_{2}^{2}}$.
If we have only discrete samples, then we need a discrete analog to Equation 4; we can set up a square system to uniquely determine $N$ coefficients:
(5) $\begin{bmatrix}\xi_{0,0}&\xi_{1,0}&\cdots&\xi_{N-1,0}\\ \xi_{0,1}&\xi_{1,1}&\cdots&\xi_{N-1,1}\\ \vdots&\vdots&\ddots&\vdots\\ \xi_{0,N-1}&\xi_{1,N-1}&\cdots&\xi_{N-1,N-1}\end{bmatrix}\begin{bmatrix}c_{0}\\ c_{1}\\ \vdots\\ c_{N-1}\end{bmatrix}=\begin{bmatrix}y_{0}\\ y_{1}\\ \vdots\\ y_{N-1}\end{bmatrix}$
where $\xi_{k,n}$ is the value of the $k^{\text{th}}$ basis function at the $n^{\text{th}}$ sample point, $x_{n}$. This equation specifies that each point $y_{n}$ should be approximated by a weighted sum of the basis functions at the corresponding $x_{n}$, so the reconstruction recipe, Equation 3, still applies. This allows us to represent the function’s value over the entire domain despite starting with only discrete samples.
The linear inverse problem in Equation 5 works for orthogonal and merely linearly independent basis sets alike, because both make the square matrix full rank. This gives the system a unique solution, although its condition number and consequent sensitivity to inexactness in the target, $\{y_{n}\}$, may vary. In the general case, solving the system requires $O(N^{3})$ operations, which is an expensive computation. However, the Fourier (section 4.2.1) and Chebyshev (section 4.2.2) bases enable much faster solutions on the order of $O(N\log N)$.
Because this method exploits all available information (all samples) to fit all basis functions, it is called global and is able to converge to a true underlying smooth function super-algebraically, i.e. faster than any $O(N^{-m})$ for $m\in\mathbb{R}$ [kutzbook, trefethen2000spectral]. This is sometimes called “infinite order accuracy” and written $O(N^{-\infty})$ [^3] If $y$ is analytic, convergence is even faster, exponential, i.e. $O(c^{N})$ for $c\in(0,1)$ [trefethen2000spectral], but in this case one should probably just use AutoDiff (section 3). [trefethen2000spectral]. This implies phenomenal accuracy, but for nonsmooth functions there is no such guarantee, limiting the applicability of spectral interpolants.
4.2.1 Differentiation with the Fourier Basis
The most ubiquitous [oppenheim, hamming] basis is the complex exponentials $e^{i\omega x}=\cos(\omega x)+i\sin(\omega x)$, where $i$ is the imaginary unit and $\omega$ is angular frequency. For real-valued functions and signals, each complex exponential is multiplied by a complex coefficient that perfectly compensates for the imaginary component to produce a real, phase-shifted sinusoid. Fitting against and reconstructing with this orthogonal basis in continuous variables gives rise to the Fourier transform, often denoted with a hat $\hat{\ }$, and its inverse:
(6) $\hat{y}(\omega)$ $=\int\limits_{-\infty}^{\infty}y(x)e^{-i\omega x}dx=\mathcal{F}\{y(x)\}$
$y(x)$ $=\frac{1}{2\pi}\int\limits_{-\infty}^{\infty}\hat{y}(\omega)e^{i\omega x}d\omega=\mathcal{F}^{-1}\{\hat{y}(\omega)\}$
If $y$ is absolutely integrable, as it must be unless it contains infinite energy, we can apply integration by parts and find a simple expression for the derivative in the transformed domain:
(7) $\mathcal{F}\{\frac{d}{dx}y(x)\}$ $=\!\int\limits_{-\infty}^{\infty}\!\underbrace{e^{-i\omega x}}_{u}\underbrace{\frac{dy}{dx}dx}_{dv}=\underbrace{y(x)e^{-i\omega x}\Big|_{-\infty}^{\infty}}_{\mathclap{\text{\shortstack{0 for Lebesgue-\\ integrable functions}}}}-\int\limits_{-\infty}^{\infty}\!\underbrace{y(x)}_{v}\underbrace{(-i\omega)e^{-i\omega x}dx}_{du}$
$=i\omega\cdot\hat{y}(\omega)$
That is, we can accomplish differentiation in the Fourier domain with mere multiplication, and we can even apply the above recursively to get a formula for the $\nu^{\text{th}}$ derivative:
(8) $y^{(\nu)}={\cal F}^{-1}\{(i\omega)^{\nu}{\cal F}\{y\}\}$
But we are performing numerical computation on sampled functions, so we must use the Discrete Fourier Transform pair:
(9) $\text{DFT: \ \ }Y_{k}$ $=\sum_{n=0}^{N-1}y_{n}e^{-i\frac{2\pi}{N}nk}$
$\text{DFT}^{-1}\text{: \ \ }y_{n}$ $=\frac{1}{N}\sum_{k=0}^{N-1}Y_{k}e^{i\frac{2\pi}{N}nk}$
where $k$ in this context is often called “wavenumber” and the coefficient $c_{k}$ from last section is renamed $Y_{k}$. The DFT and its inverse can be computed in $O(N\log N)$ by the Fast Fourier Transform (FFT) algorithm [cooley1965algorithm], allowing us to avoid inverting the matrix in Equation 5.
There is a discrete analog of Equations 7 and 8 as well [johnson-fft-notes, spectral-derivatives], which essentially band-limits the spectral reconstruction to frequencies of lower absolute value:
(10) $Y^{(\nu)}_{k}=\begin{cases}(ik)^{\nu}\cdot Y_{k}&k<\frac{N}{2}\\ (i\frac{N}{2})^{\nu}\cdot Y_{k}&k=\frac{N}{2}\text{ and }\nu\text{ even}\\ 0&k=\frac{N}{2}\text{ and }\nu\text{ odd}\\ (i(k-N))^{\nu}\cdot Y_{k}&k>\frac{N}{2}\end{cases}$
This gives rise to the highly accurate and fast solution procedure in 2. But there is a catch: The FFT assumes its input signal is defined on the domain $x\in[0,2\pi)$ and furthermore requires that the signal in question is periodic on this interval, because $e^{-i\frac{2\pi}{N}n(k+N)}=e^{-i\frac{2\pi}{N}nk}$ and $e^{i\frac{2\pi}{N}(n+N)k}=e^{i\frac{2\pi}{N}nk}$ in Equation 9, and therefore $Y_{k+N}=Y_{k}$ and $y_{n+N}=y_{n}$. If $y$ is periodic on some other interval of length $T$, we can map $x\in[x_{0},x_{0}+T)$ to $\theta\in[0,2\pi)$, find the scaled $\nu^{\text{th}}$ derivative, and rescale by $(\frac{2\pi}{T})^{\nu}$ without a problem. But if $y$ is aperiodic, then when the DFT implicitly periodically extends $y$, there will be discontinuities or corners in the extended function, and these are impossible to fit with a finite number of smooth, sinusoidal basis functions, resulting in artifacts known as Gibbs phenomenon, illustrated in Figure 7.
Algorithm 2 Differentiation via FFT
1:Find the Fourier coefficients $Y_{k}=\text{FFT}(y_{n})$.
2:Multiply by appropriate $(ik)^{\nu}$, 0, or $(i(k-N))^{\nu}$ from Equation 10 to make Fourier coefficients of the derivative, $Y^{(\nu)}_{k}$.
3:Inverse transform to obtain samples of the derivative function $y^{(\nu)}_{n}=\text{FFT}^{-1}(Y^{(\nu)}_{k})$
Figure 7: From [kutzbook], periodic extension of a ramp function results in discontinuities at the domain endpoints, which cannot be perfectly reconstructed with 8, 16, or any number of smooth sines and cosines, leading to overshoots and wobbles known as Gibbs phenomenon.
4.2.2 Differentiation with a Polynomial Basis: Chebyshev
Polynomial bases have the requisite expressive power to represent aperiodic functions and come in many flavors, each with its own properties: Legendre polynomials are natively orthogonal on [1,1][-1,1], while Chebyshev polynomials are orthogonal only with the inclusion of the weight function $1/\sqrt{1-x^{2}}$ in the inner product integral over [1,1][-1,1]. Hermite polynomials are orthogonal on $\infty$$\infty$] with weight function $e^{-x^{2}}$ [dedalus], but Bernstein polynomials, defined on $[0,1]$, are not orthogonal [spectral-derivatives], just linearly independent. Each has its own optimal sampling grid [dedalus] (Figure 8) to best avert Runge’s phenomenon (Figure 9) by lowering the condition number of the matrix in Equation 5.
Figure 8: From [dedalus], optimal sampling grids across interval domains, also called “collocation points”, for a few common choices of basis. Each basis is defined on its own canonical interval, e.g. $0,2\pi)$ for Fourier and [1,1][-1,1] for Chebyshev, so other intervals are affine-transformed before basis function fit, which is easily reversible to return to the original domain.
Figure 9: From [kutzbook]. With equispaced sampling grids, as in (a) and (b), high-order polynomial fits exhibit wobbles at the edges of the domain, known as Runge’s phenomenon. This can be mitigated by sampling at nodes better suited to the polynomial basis, such as the Chebyshev-Lobatto nodes in the case of Chebyshev polynomials, shown in (c) and (d).
Of the polynomial bases, the Chebyshev basis is the most practically useful, because it is identical to a basis of cosines under the change of variables $x=\cos(\theta),\ x\in[-1,1],\ \theta\in[0,\pi]$ per Equation 11, depicted in Figure 10 [trefethen2000spectral, dedalus].
(11) $T_{k}(\cos\theta)=\cos(k\theta)$
$y(x)=\sum_{k=0}^{N-1}a_{k}T_{k}(x)\ ;\quad y(\theta)=\sum_{k=0}^{N-1}a_{k}\cos(k\theta)$
Figure 10: From [dedalus], relationship of Chebyshev domain and Fourier Domain. Chebyshev polynomials look like the xyxy-plane shadows of cosines wrapped around a cylinder. Notice the cosines are horizontally flipped.
Coefficients of a cosine basis can be found with the Discrete Cosine Transform (DCT), [^4] Be cautious that, using discrete $\theta_{n}=\frac{\pi n}{N-1}$, the DCT-I-1 says $y_{n}=\frac{1}{2(N-1)}\big(Y_{0}+Y_{N-1}(-1)^{n}+2\sum_{k=1}^{N-2}Y_{k}\cos(\frac{\pi kn}{N-1})\big)$, but in the Chebyshev expansion $y_{n}=\sum_{k=0}^{N-1}a_{k}\cos(\frac{\pi kn}{N-1})$. $Y_{0}$ is multiplied by $1=\cos(\frac{\pi\cdot 0\cdot n}{N-1})$ and $Y_{N-1}$ by $(-1)^{n}=\cos(\frac{\pi\cdot N-1\cdot n}{N-1})$, so the first and last terms can be put inside the sum but are twice as large as the others, and all $Y_{k}$ have to be scaled to produce $a_{k}$ [spectral-derivatives]. based on a modification of the FFT, in $O(N\log N)$, so by reduction to this case, Chebyshev coefficients can be found equally fast. Recall the FFT requires equispaced points in variable $\theta\in[0,2\pi)$, as does the DCT with $\theta\in[0,\pi]$, where half the interval is dropped because the function is assumed to be even, making information from the latter half redundant. Thus, to take advantage of these algorithms, samples in $x$ need to come from $x_{n}=\cos(\theta_{n})=\cos(\frac{\pi n}{N-1})\in[-1,1]$ [trefethen2000spectral], which are the Chebyshev-Lobatto nodes [brown-cheb], corresponding to the $N$ extrema of $T_{N-1}$ [dedalus].
Once the Chebyshev transform is achieved, then, analogous to the power rule for ordinary representations of polynomials as power series, we can find Chebyshev series coefficients of the derivative in $O(N)$ using a recurrence relation [brown-cheb], which can be manipulated to produce Equation 12 [spectral-derivatives], implemented in numpy as chebder [numpy-chebder]:
(12) $\frac{d}{dx}T_{k}(x)=k\cdot\begin{cases}2\sum_{\text{odd }j>0}^{k-1}T_{j}(x)&\text{for even }k\\ -1+2\sum_{\text{even }j\geq 0}^{k-1}T_{j}(x)&\text{for odd }k\end{cases}$
Algorithm 3 Differentiation with the Chebyshev Basis
1:Find the Chebyshev coefficients with $a_{k}\propto Y_{k}=\text{DCT}(y_{n})$
2:Use the series recurrence, Equation 12, to make Chebyshev coefficients of the derivative, $a^{(\nu)}_{k}$.
3:Scale to make DCT coefficients, $Y^{(\nu)}_{k}$, and inverse transform to obtain samples of the derivative function $y^{(\nu)}_{n}=\text{DCT}^{-1}(Y^{(\nu)}_{k})$
4.2.3 Recommendations
Fourier spectral differentiation is best used with periodic signals. By including higher frequencies, the fit can get closer at discontinuities and corners, but Gibbs phenomenon will always remain [oppenheim, trefethen2000spectral]. In some cases this problem can be alleviated or mitigated by concatenating an aperiodic signal with its mirror image (even extension) or negated mirror image (odd extension) to produce a truly periodic signal or something very close to periodic, as in Figure 11. The FFT is so fast that transforming a double-length signal is only marginally slower. But this trick does not produce a perfectly periodic extension for general signals.
Figure 11: Examples of aperiodic functions that can be periodically extended, $\sin(\frac{3\theta}{2})$ on $\theta\in[0,2\pi)$ and $\frac{1}{1+e^{-\theta}}$ on $\theta\in[-5,5)$. Black + marks indicate function samples, and colorful + marks show the Fourier-based differentiation algorithm’s answers for subsequent derivatives. Answers remain essentially perfect for the truly periodic extension but start to drift for higher derivatives of the only-nearly-periodic extension.
3 offers a method with exceptional computational efficiency and accuracy while also being unconstrained by boundary conditions. However, function samples need to come from special locations or an affine transform thereof, and the method should only be used in the absence of noise (see section 5, particularly Figure 19). The Matlab package chebfun [driscoll2014chebfun] and Python packages spectral-derivatives [spectral-derivatives] and dedalus [dedalus] exploit Chebyshev polynomials to allow for accurate and efficient derivative estimates of arbitrary smooth functions.
4.3 Finite Elements
 [^5] There is also Finite Volumes, which emphasizes conservation of fluxes across boundaries rather than fit to a basis, introducing integrals that resemble the weak form of Finite Elements (section 4.3.1). In fact, for conservation-law PDEs in divergence form, FVM can be viewed as a Galerkin method (see below) with piecewise constant test functions, though it is conventionally regarded as a distinct scheme.Instead of projecting a function onto basis functions that span the whole domain, like Spectral Methods (section 4.2), consider basis functions with local, overlapping domains, as shown in Figure 12 and Figure 13.
Figure 12: From [comsol], Finite Element basis functions, $\psi$, with only local support, of type Lagrange polynomial with order 1. (a) Equispaced, (b) on a mesh that captures a steeper section with higher fidelity, (c) in 2D. In (a) and (b) a target function is shown in blue, with the best-fit sum of basis functions shown as a dotted red line.
[…]
Figure 13: From [langtangen], Finite Element basis functions, this time called $\varphi$, of type Lagrange polynomial, (a) linear, (b) quadratic, (c) cubic. Boundaries between elements are shown as pink dotted lines. As the order increases, the number of basis functions associated with each element increases, as does the number of “nodes” circled in red where all-but-one of these basis functions is zero. Within an element, the basis forms a partition of unity, i.e. the basis functions sum to 1.
Basis functions of this kind are capable of representing a certain family of functions, together forming the “trial space”, often referred to in the literature [langtangen, larson, dokken] by the symbol $V_{h}$:
$V_{h}=\{v:v\in C^{d}(\Omega),\ v|_{\Omega_{e}}\in\text{span}(\{\varphi_{k}\})=\text{the basis space}\}$
That is, the space is the set of functions $v$ that satisfy: (1) $d$ times continuous differentiability over domain $\Omega$ [^6] We use $\Omega$ to suggest that the domain and function can be multidimensional, but we primarily stay in 1D for illustration. and (2) can be built piecewise by stringing together elements $v|_{\Omega_{e}}$, which are composed from the basis functions and only nonzero over subdomains $\Omega_{e}$ of size related to a scalar parameter $h$ (hence the subscript). Often, the basis space is Lagrange polynomials of low degree, because these are cheap to compute. If we demand only that the resulting $v$ be continuous, then $d=0$, and we build from piecewise linear components, resulting in the faceted appearance of many Finite Elements solutions. Using high-order basis functions is sometimes referred to as “Spectral Elements”, because each element’s fit is comprised of many basis functions, reminiscent of Spectral Methods.
Building from piecewise elements has great benefits for flexible representation of complicated functions on irregular domains, such as shown in Figure 14. Whereas ordinary Finite Difference requires a regular grid, [^7] Technically, the Finite Difference equations in section 4.1 can be extended to the irregular sampling case, as described in section 8.1, but this is uncommon. and Spectral Methods work most naturally on rectangular domains without holes, elements are suited to work in a variety of shapes, including line segments in 1D, triangles and squares in 2D, and tetrahedra and prisms in 3D.
Figure 14: From [huebner], a comparison of domain discretization for (a) Finite Difference and (b) Finite Elements.
But a piecewise function may not have the expressive power to fully represent the target function, which typically lives in a more general function space, often called $V$, for instance:
$V=\{v:\|v\|_{L_{2}(\Omega)}<\infty,\|\nabla v\|_{L_{2}(\Omega)}<\infty,v(\partial\Omega)=\text{boundary conditions}\}$
Here the $L_{2}$ norm of $v$, i.e. the square root of the inner product of $v$ with itself, which is an integral over the domain, has to converge, as does the norm of $v$’s first derivative ($\nabla$ if there are multiple dimensions), and $\partial\Omega$ is the domain boundary.
Our first task is to represent $y\in V$ with $y_{h}\in V_{h}\subset V$, as illustrated in Figure 15. We can naturally measure distance between these objects with the induced $L_{2}$ norm:
(13) $\|y-y_{h}\|_{L_{2}(\Omega)}^{2}=\int\limits_{\Omega}(y-y_{h})\overline{(y-y_{h})}d\Omega$
where the overbar means complex conjugate. The minimizer of this equation is provably [^8] See the short proof of theorem 1.1 in [larson]. $y_{h}=P_{h}y$, the projection of $y$ onto the subspace $V_{h}$. But we cannot minimize this equation directly, because continuous functions are infinite-dimensional: Fitting at all continuum points is impossible to compute.
Figure 15: From [larson]. (a) Notional illustration of a function, here called $f$, being $L_{2}$-projected onto the space $V_{h}$. (b) $f(x)=2x\sin(2\pi x)+3$ and its $L_{2}$-projection $P_{h}f(x)$ using the mesh $[0,0.2,0.4,0.6,0.8,1]$. The fit does not perfectly fall on the function at the mesh points, the boundaries of elements, because the fit is not linear interpolation; rather it is best fit in an $L_{2}$-distance, inner-product induced-norm sense. See Equation 13
Instead we can use what is called a Galerkin method, one among a group of methods known as the weighted residuals approach [huebner], to turn the problem into a discrete set of constraints and solve these to obtain $y_{h}$. The trick is to notice that $y-y_{h}$ (called $f-P_{h}f$ in Figure 15(a)) is orthogonal to all $v\in V_{h}$, which in this context are called “test functions”:
$\langle y-y_{h},v\rangle=0\Leftrightarrow\int\limits_{\Omega}(y-y_{h})\overline{v}\ d\Omega=0$
Since $y_{h}$ is composed of basis functions, we can write it as $y_{h}=\sum_{j}c_{j}\varphi_{j}(x)$, where the $\{c_{j}\}$ are coefficients. The total number of basis functions is related to the mesh size, $N$, and the order of the basis, which can drive up the number of nodes per element. For linear basis functions as in Figure 12(a), elements only have nodes at their boundaries, so $j\in\{0,...,N-1\}$.
A Galerkin projection uses the same family of functions for the basis and test sets. We have actually already encountered something which can be thought of as Galerkin projection, the Fourier transform, or more specifically the Fourier series [oppenheim, fourier-series], which maps from a continuous function on a periodic interval, say $\pi,\pi]$, to discrete coefficients, using complex exponentials:
(14) $\int\limits_{-\pi}^{\pi}(y(x)-y_{h}(x))e^{-ikx}dx=0$
The reconstruction sum $y_{h}(x)=\sum_{j=0}^{N-1}c_{j}e^{ijx}$ leaves us with $N$ unknown coefficients, so we need $N$ equations to lock them down. These we can obtain through Equation 14 by projecting against test functions with different $k$. Plugging in the ansatz for $y_{h}$, we find:
(15) $\rightarrow\int\limits_{-\pi}^{\pi}y(x)e^{-ikx}dx=\int\limits_{-\pi}^{\pi}\Big(\sum_{j=0}^{N-1}c_{j}e^{ijx}\Big)e^{-ikx}dx\quad\text{for }k=0,.,N-1$
$\rightarrow\underbrace{\langle y(x),e^{ikx}\rangle}_{\text{\small$b_{k}$}}=\sum_{j=0}^{N-1}c_{j}\underbrace{\int\limits_{-\pi}^{\pi}e^{ijx}e^{-ikx}dx}_{\text{\footnotesize$\begin{cases}2\pi&j=k\\ 0&j\neq k\end{cases}$}}=2\pi c_{k}\quad\text{for }k=0,.,N-1$
$\rightarrow\mathbb{I}\mathbf{c}=\frac{1}{2\pi}\mathbf{b}$
The coefficients in c are of course the Fourier coefficients.
A linear inverse problem of this kind is typical in Finite Elements. In general, the matrix multiplying the coefficients, often called the “mass matrix”, $\mathbf{M}$, is not identity, but thankfully using local, piecewise basis functions causes the entries of $\mathbf{M}$ to be sparse, because most pairs of basis functions do not overlap, so their products integrate to 0. This allows the system to be solved more efficiently than by naive $O(N^{3})$ matrix inversion.
The approximation error of fitting the function $y$ with piecewise $y_{h}$ is $O(h^{\alpha})$, where $h$ is related to the side-length of a typical element [comsol, larson]. This is an algebraic bound, similar to that of Finite Difference, and indeed Finite Elements can be thought of as an alternative form of Finite Difference that interpolates over patches of the domain [kutzbook, langtangen]. The parameter $\alpha$ can be difficult or impossible to estimate for all but simple problems a priori, so software solvers often use their solutions to calculate a posteriori error estimates and choose how to make refinements [comsol, langtangen].
4.3.1 Differentiation with Finite Elements
We have now reached a curious juncture: We can compute a piecewise basis-function representation of a function we wish to differentiate, but to do so involves inner product integrals. In practice, on a computer integrals are found via quadrature rules [larson], which use samples from particular “quadrature points” to compute the integral of a polynomial interpolant. If we provide the method with samples $y_{n}$ rather than function $y$, and cannot guarantee the spacing of these points, then trapezoidal rule (or its higher-dimensional analog) is the only integration rule robust to this scenario, requiring only quadrature points at element boundaries. But the trapezoidal rule interpolates between vertices linearly, thereby assuming that $y\in V_{h}$ already.
So the first important thing to recognize is Finite Element methods are best for working with analytic governing equations, not data samples. Indeed, choosing where to sample (the irregular mesh) is part of the setup, so the method needs to be able to sample initial conditions, $y_{0}$, at arbitrary points to get started. However, unlike the Automatic Differentiation case (section 3), we expect the exact solution $y$ not to be analytic as the system evolves. Technically, piecewise $y_{h}$ remains analytic, but it is complicated, much friendlier to compute than to write down.
Second, since $y_{h}$ is intentionally piecewise, it will have corners and kinks at element boundaries, making it not strictly differentiable at those locations, like the functions in Figure 16. So instead of directly taking derivatives to solve a governing equation, the solution is found in the “weak” (integral) form, which only requires that the inner products of the two sides of the equation with test functions be equal, thereby allowing discontinuities in the first derivative [comsol].
1-11111xxReLU1-11111xxHeaviside1-111xxDirac Delta (symbolic)
Figure 16: Examples of functions where the notion of a weak derivative is useful. The Rectified Linear Unit (ReLU) has a sharp corner at $x=0$, making it not strictly differentiable. But intuitively its derivative should look something like a Heaviside function, which in turn has a weak derivative given by the Dirac Delta. This notion is made rigorous by the machinery of “generalized” functions (also called “distributions”) [weak-derivs], members of which can be written as a functional which maps from function to value via an integral against a test function, essentially identical to the form in Galerkin projection.
This setup matches the Galerkin method naturally, and it comes with the added benefit, thanks to an integration by parts trick [^9] The same manipulation is famously used in Calculus of Variations to transfer a differential operator from the variation to the Lagrangian while spawning a boundary term, thereby enabling a factorization that yields the Euler-Lagrange equation. In higher dimension, this trick may be called Green’s First Identity [langtangen], a special case of Divergence Theorem. (demonstrated below), that the solution need only be (weakly) differentiable one fewer times in space than in the original PDE, thereby enabling us to describe it with a lower-order basis, which means fewer nodes (see Figure 13) and a smaller system of equations, especially helpful for saving computation in higher dimensions. [^10] The weak formulation may not be strong, but it is powerful.
It is instructive to see an example. In the following we use subscript $x$ and $t$ as a shorthand to mean partial derivatives. The weak solution for the 1D wave equation, often named $u$ as in section 3.1 rather than $y$, can be formed as:
$u_{tt}=u_{xx},\quad x\in[0,L]$
$u(x,0)=u_{0}(x),\quad u_{t}(x,0)=0,\quad u(0,t)=u(L,t)=0$
[…]
Using $u_{xx}dx=d(\frac{d}{dx}u)$, we can integrate the right-hand side by parts:
$\int\limits_{0}^{L}v\ du_{x}=u_{x}v\Big|_{0}^{L}-\int\limits_{0}^{L}u_{x}dv$
If we require the test functions $v$ to be 0 at the edges of the domain, then the first term disappears. We can additionally use $dv=\frac{d}{dx}v\ dx$ to obtain:
$\int\limits_{0}^{L}u_{tt}v\ dx=-\int\limits_{0}^{L}u_{x}v_{x}\ dx$
[…]
$\int\limits_{0}^{L}\sum_{j=1}^{N-2}\ddot{c}_{j}(t)\varphi_{j}(x)\varphi_{k}(x)\ dx=-\int\limits_{0}^{L}\sum_{j=1}^{N-2}c_{j}(t)\varphi^{\prime}_{j}(x)\varphi^{\prime}_{k}(x)\ dx\quad\text{for }k=1,...,N-2$
$\rightarrow$ $\sum_{j=1}^{N-2}\ddot{c}_{j}(t)\underbrace{\int\limits_{0}^{L}\varphi_{j}(x)\varphi_{k}(x)\ dx}_{M_{jk}}=-\sum_{j=1}^{N-2}c_{j}(t)\underbrace{\int\limits_{0}^{L}\varphi^{\prime}_{j}(x)\varphi^{\prime}_{k}(x)\ dx}_{K_{jk}}\quad\text{for }k=1,...,N-2$
$\rightarrow\mathbf{M}\ddot{\mathbf{c}}_{1:N-2}(t)+\mathbf{K}\mathbf{c}_{1:N-2}(t)=0$
The $(N-2)\!\times\!(N-2)$ entries of $\mathbf{M}$ and $\mathbf{K}$ do not change over time. We have transformed the PDE into a system of ODEs. Initial coefficients c can be found from initial conditions and then forward-solved with a method of lines scheme and steppers like Runge-Kutta. If the PDE instead describes a steady state with unknown $u$, the derivation culminates in an algebraic equation rather than ODEs, similar to the Fourier series example in Equation 15. For a further example, see the derivation for the 2D Poisson equation in any of [langtangen, larson, dokken, olver, uiuc-galerkin], which is also instructive for generalizing to more spatial dimensions.
The overall solution process of Finite Elements is quite involved, from mesh generation, to boundary condition enforcement, to linear inverse problem (or even nonlinear problem [langtangen, larson]) setup, to numerical solution methods (e.g. Newton-Raphson), to error estimation and refinement. Thankfully, most of these steps can be automated, which is exactly what is done by commercial packages like COMSOL and Ansys and open-source ones like FEniCSx [fenicsx].
4.3.2 Recommendations
Finite Elements is not well suited to differentiating a typical array of numbers representing a signal, but it is the most versatile method for solving governing PDEs on arbitrary domains with arbitrary boundary conditions, which are common in engineering, especially in two and three dimensions. Due to the use of piecewise functions and the weak form, Finite Elements are also capable of representing true discontinuities like shockwaves, [^11] It is possible to approximate a discontinuity with Spectral Methods, as demonstrated by Dedalus [dedalus], but the solution will require a high number of modes to represent and is still technically smooth, with rounded corners and a steep connecting segment. which can be present in the solutions of hyperbolic PDEs [olver]. Due to its complexity, Finite Elements is best done via established software. Commercial packages require a license, while the free, open-source FEniCSx requires formulating a PDE in the weak form [dokken].
5 Noise
In practice, most real-world data is corrupted due to imperfect sensing (“measurement noise”), unknown forcing in the dynamics (“process noise”), or statistical or numerical artifacts (such as sampling noise or quantization noise). We model this by considering observations to be the additive result of some unknown true signal and noise, $\eta$:
(16) $y=y_{\text{true}}+\eta$
Noise can be deadly for sample-based differentiation methods, because small deviations in sample values can lead to large deviations of slope, especially when considering a constant measurement noise level with shrinking $\Delta x$. This effect is only magnified for calculations of curvature and higher-order derivatives, $\frac{d^{\nu}y}{dx^{\nu}}$.
However, if we have reason to believe the data is generated by an underlying process that evolves smoothly along our independent variable of interest (whether time or space), and we sample sufficiently densely to capture these variations, then meaningful changes in the signal can be represented by a sum of relatively low-wavenumber Fourier basis functions. This is a central observation in Signal Processing: For a naturally-occurring signal, “energy” in the sense of Parseval’s Relation [oppenheim], Equation 17, must be finite, and consequently its spectrum must fall off toward zero as frequency increases [hamming].
(17) $\int\limits_{-\infty}^{\infty}|y(x)|^{2}dx=\frac{1}{2\pi}\!\!\!\int\limits_{-\infty}^{\infty}|\hat{y}(\omega)|^{2}d\omega$
It is worth noting that energy also empirically clusters in lower modes when signals are decomposed using other functions, e.g. polynomials (section 4.2.2) or neural nets [driscoll2014chebfun, spectral-bias-nn]. This is even more so the case when a basis is tailored to represent a particular signal, like Singular Value Decomposition (SVD) [ddse], as demonstrated in Figure 17.
Figure 17: We can stack red, blue, and green channels of an image, perform SVD on the 2D matrix, and reshape $U[:,:r]\cdot\Sigma[:r]\cdot V^{T}[:r]$ to produce reconstructions with $r$ modes. In this case, the first 10 modes alone contain more than half the signal energy. Source image from Smithsonian.
Noise must obey this same energy decay, but its Fourier spectrum tends to have less bias toward low frequencies, because noise contains sharp variations that can only be modeled with higher frequencies. Alternative bases tend not to separate high-frequency noise as consistently across the domain [spectral-derivatives]. Figure 18 sketches the spectrum of a noisy signal, which can be thought of as the sum of signal and noise spectra, by consequence of Equation 16 and the linearity of the Fourier transform. It also depicts the phenomenon of aliasing, where a sampled higher frequency looks like and adds to a lower one, because it takes >2>\!2 samples per cycle to truly resolve a sinusoid [oppenheim], while a spectral reconstruction is composed of lower-magnitude wavenumbers (see section 4.2.1).
[…]
Figure 18: Reproduced from [spectral-derivatives], typical energy spectrum in the frequency (i.e. Fourier) domain of a noisy signal, before and after sampling. $f_{s}$ is a sampling rate of our choosing. Equispaced samples of a sinusoid with frequency just over the Nyquist limit, $\frac{1}{2}f_{s}+\epsilon$, can be fit equally well by a single sinusoid of frequency just under the Nyquist limit, $\frac{1}{2}f_{s}-\epsilon$, a phenomenon known as “aliasing” due to this 1:1 relationship. This causes a folding pattern in the spectrum, here shown explicitly, although the final spectrum of a sampled signal is the sum of these folds up to infinite frequency. The infinite sum must converge, so the signal contains ever-decreasing power at higher frequencies. With increasing $f_{s}$, the folding noise spectrum’s energy is wider-spread, causing less overlap with the signal.
The band-separability of signal from noise illustrated in Figure 18 is the spiritual core of noise reduction techniques: At bottom, all accomplish some kind of low-pass filtering. The following sections explore myriad such approaches and present a strategy for choosing algorithm hyperparameters based on the frequency content of measured data.
5.1 Recommendations
When a signal is known to be periodic, applying an ideal low-pass filter, i.e. 2 with higher modes zeroed out before FFT-1, separates signal from high-frequency noise efficiently and optimally—although some noise spectrum will still unavoidably overlap the signal’s spectrum. Beware, one should not attempt to use the Chebyshev basis in the presence of noise, because polynomials have nonuniform oscillations and therefore do not filter noise equally well throughout the domain. Even worse, noise-induced fit inaccuracies propagate and compound from higher modes to lower ones during differentiation due to Chebyshev mode coupling [^12] By contrast with the aliasing of sinusoids, the best fit to samples from a single higher-degree polynomial requires many different lower-degree polynomials. [brown-cheb], resulting in systematic error toward the domain edges, as depicted in Figure 19. Likewise, the eigenmodes of a signal, ideal for parsimonious representation [ddse], are generally not best for separating noise due to their nonuniform frequency content across the domain. So in practice, with noisy, aperiodic data, one of the many methods discussed in the following two sections is likely to provide the most accurate results.
Figure 19: From [spectral-derivatives], the function $y(x)=e^{x}\sin(5x)$ is sampled with noise $\sim\mathcal{N}(\mu=0,\sigma=0.1)$ at 100 points and fit with 100 Chebyshev polynomials, only the first 20 of which are kept after a filtering step. The fit is characteristically prone to overfit near domain edges where the basis functions are steeper and have more wobbles, leading to dramatic errors in derivative estimates. Raising the cutoff makes this phenomenon worse, while lowering the cutoff leads to poor fits by degrading the method’s expressive power. This phenomenon is not easily sidestepped by padding, because cosine-spaced samples cannot be padded without breaking cosine-spacing and $O(N\log N)$ runtime.
6 Differentiating Noisy Data Using Prior Knowledge
In some cases, additional information may be available, such as a model of the system that produced data, a model of noise, or relevant synchronized data streams. Incorporating this knowledge can help distinguish signal from noise. The most dominant framework for this situation is Kalman filtering/smoothing, [^13] The filtering problem is to make estimates online, in one direction, whereas smoothing assumes a whole data series is available offline, enabling a backwards pass. See section 6.3. which models the true signal, often along with one or more of its derivatives, as a hidden state and calculates estimates by weighted sum of noisy measurements and dynamics-based predictions.
The standard Kalman filter handles only linear dynamics, uses a mean squared error (MSE) penalty metric, and is the maximum a priori (MAP) estimator (section 6.2) only for zero-mean Gaussian white noise disturbances. This is fairly limiting and reflects the original formulation’s place in history, [^14] The Kalman filter was meant to generalize the Wiener filter for non-stationary processes [mit-wiener-kalman], although Wiener can explicitly handle colored noise (correlated in time) while Kalman requires modification with shaping filters and augmented state to turn such noise white [crassidis_junkins]. so we contextualize by surveying a variety of generalizations, including the Luenberger observer [luenberger] (section 6.1); robust variants designed to handle more exotic noise distributions, outliers, sparsity, and constraints on the state [aravkin] (section 6.4); $H$ filtering [robust-control] (Hardy space); and a modification for use with nonlinear systems [UKF] (section 6.5).
Common symbols used in this framework differ from the notation used so far, so we give them in Table 3.
Table 3: Common symbols used with Kalman Filters.
$\mathbf{x}$ true signal (or state), not directly observable
$\mathbf{\hat{x}}$ estimated signal (state)
$\mathbf{e}$ state estimation error, $\mathbf{x}-\mathbf{\hat{x}}$
$\mathbf{y}$ noisy measurement
$\mathbf{u}$ known control input or synchronized data stream
$\mathbf{A}$ sometimes also called $\mathbf{F}$ or $\boldsymbol{\Phi}$, linear state evolution matrix or Jacobian of nonlinear state evolution function $\boldsymbol{f}$
$\mathbf{B}$ linear control matrix or relation of known states to hidden states
$\mathbf{C}$ sometimes also called $\mathbf{H}$ or $\boldsymbol{\Phi}$ (confusingly), linear measurement matrix or Jacobian of nonlinear measurement function $\boldsymbol{h}$
$\mathbf{w}$ process noise
$\mathbf{v}$ measurement noise
$\mathbf{Q}$ process noise covariance
$\mathbf{R}$ measurement noise covariance
$\mathbf{P}$ covariance of state estimate error
$\mathbf{K}$ the Kalman gain matrix, filter parameters
$p$ a probability density function
$\mathcal{N}$ the normal distribution, univariate or multivariate, parameterized by scalar or vector mean, $\mu$ or $\boldsymbol{\mu}$, and standard deviation or covariance, $\sigma$ or $\boldsymbol{\Sigma}$
6.1 The Classic Kalman Filter
The Kalman filter solves the minimum mean squared error optimization problem, for a discrete update from index n1n-1 to $n$:
(18) $\underset{\mathbf{K}}{\text{minimize}}$ $\mathbb{E}[\|\mathbf{x}_{n}-\mathbf{\hat{x}}_{n}\|_{2}^{2}]$
s.t.(i) $\mathbf{x}_{n}=\mathbf{A}\mathbf{x}_{n-1}+\mathbf{B}\mathbf{u}_{n}+\mathbf{w}_{n},\ \mathbf{w}_{n}\sim\mathcal{N}(0,\mathbf{Q})$
(ii) $\mathbf{y}_{n}=\mathbf{C}\mathbf{x}_{n}+\mathbf{v}_{n},\ \mathbf{v}_{n}\sim\mathcal{N}(0,\mathbf{R})$
(iii) $\mathbf{\hat{x}}_{n}=\mathbf{A}\mathbf{\hat{x}}_{n-1}+\mathbf{B}\mathbf{u}_{n}+\mathbf{K}(\mathbf{y}_{n}-\mathbf{\hat{y}}_{n})$
(iv) $\mathbf{\hat{y}}_{n}=\mathbf{C}(\mathbf{A}\mathbf{\hat{x}}_{n-1}+\mathbf{B}\mathbf{u}_{n})$
That is, assuming linear dynamics and measurement models, perturbed by noise drawn from zero-mean Gaussian probability distributions ((i) and (ii)), form the state estimate by weighted sum of (1) a prediction from the dynamics and (2) the consequent measurement error ((iii) and (iv)). Then find the weight matrix projecting from measurement space to state space that minimizes the Mean Squared Error (MSE), the expected value of the $\ell_{2}$ norm of the state estimate error.
Noise terms in the model are characterized by vector means (0 for both) and covariance matrices, $\mathbf{Q}=\text{cov}(\mathbf{w},\mathbf{w})$ and $\mathbf{R}=\text{cov}(\mathbf{v},\mathbf{v})$, because states and measurements may have multiple variables. Covariance of a random vector $\boldsymbol{\chi}$ with itself is given as:
$\text{cov}(\raisebox{0.95pt}{\small$\boldsymbol{\chi}$},\raisebox{0.95pt}{\small$\boldsymbol{\chi}$})=\mathbb{E}[(\raisebox{0.95pt}{\small$\boldsymbol{\chi}$}-\mathbb{E}[\raisebox{0.95pt}{\small$\boldsymbol{\chi}$}])(\raisebox{0.95pt}{\small$\boldsymbol{\chi}$}-\mathbb{E}[\raisebox{0.95pt}{\small$\boldsymbol{\chi}$}])^{T}]=\mathbb{E}[\raisebox{0.95pt}{\small$\boldsymbol{\chi}$}\raisebox{0.95pt}{\small$\boldsymbol{\chi}$}^{T}]-\mathbb{E}[\raisebox{0.95pt}{\small$\boldsymbol{\chi}$}]\mathbb{E}[\raisebox{0.95pt}{\small$\boldsymbol{\chi}$}]^{T}$
where the last step uses the fact $\mathbb{E}[\raisebox{0.95pt}{\small$\boldsymbol{\chi}$}]$ is a constant, so the multiplicative cross terms both have the same value as the last term, but with opposing sign: $\mathbb{E}[\mathbb{E}[\raisebox{0.95pt}{\small$\boldsymbol{\chi}$}]\raisebox{0.95pt}{\small$\boldsymbol{\chi}$}^{T}]=\mathbb{E}[\raisebox{0.95pt}{\small$\boldsymbol{\chi}$}\mathbb{E}[\raisebox{0.95pt}{\small$\boldsymbol{\chi}$}]^{T}]=\mathbb{E}[\raisebox{0.95pt}{\small$\boldsymbol{\chi}$}]\mathbb{E}[\raisebox{0.95pt}{\small$\boldsymbol{\chi}$}]^{T}$. Because these covariance matrices are formed as products of vectors with their transposes, they are always symmetric and positive semidefinite and can be eigendecomposed:
[…]
where $\mathbf{V}$ is an orthonormal basis aligned with principal directions of variation, and $\boldsymbol{\Lambda}$ is filled with variances, $\sigma\geq 0$. This provides a neat way to visualize the standard deviation: Multiply points on a unit spheroid by the matrix $\mathbf{V}\boldsymbol{\Lambda}^{1/2}$, which performs scaling followed by rotation, to produce points on an ellipsoid, as in Figure 20.
$\sigma_{0}$$\vec{v}_{0}$$\vec{v}_{1}$$\sigma_{1}$
Figure 20: Visualization of one standard deviation of a 2D Gaussian distribution with mean $\boldsymbol{\mu}=0$ and covariance $\boldsymbol{\Sigma}=\mathbf{V}\boldsymbol{\Lambda}\mathbf{V}^{T}$.
An illuminating reformulation of Equation 18, purely in terms of state estimate error, $\mathbf{e}=\mathbf{x}-\mathbf{\hat{x}}$, can be found by substituting (ii) and (iv) into (iii) and subtracting from (i):
(19) $\hskip 40.00006pt\underset{\mathbf{K}}{\text{minimize}}$ $\mathbb{E}[\|\mathbf{e}_{n}\|_{2}^{2}]$
s.t. $\mathbf{e}_{n}=(\mathbb{I}-\mathbf{K}\mathbf{C})\mathbf{A}\mathbf{e}_{n-1}+\underbrace{(\mathbb{I}-\mathbf{K}\mathbf{C})\mathbf{w}_{n}-\mathbf{K}\mathbf{v}_{n}}_{\text{still Gaussian}}$
The constraint in Equation 19 gives the error dynamics. For error to decay, it is necessary that the closed-loop discrete-time map, $(\mathbb{I}-\mathbf{K}\mathbf{C})\mathbf{A}$, has eigenvalues inside the unit circle (magnitude <1\!<\!1). This is achievable by pole placement when the discrete-time pair $(\mathbf{A},\mathbf{C}\mathbf{A})$ is observable [xu-chen], giving rise to a Luenberger observer [luenberger], of which the Kalman filter is a special case that additionally accounts for noise by solving the minimization problem.
The error distribution is itself also a zero-mean Gaussian, because error is the sum of a deterministic part, that tends toward zero for a stable observer, and scaled, zero-mean Gaussians. Its mean is evaluated as $\mathbb{E}[(\mathbb{I}-\mathbf{K}\mathbf{C})(\mathbf{A}\mathbf{e}_{n-1}+\mathbf{w}_{n})-\mathbf{K}\mathbf{v}_{n}]=(\mathbb{I}-\mathbf{K}\mathbf{C})(\mathbf{A}\mathbb{E}[\mathbf{e}_{n-1}]+\mathbb{E}[\mathbf{w}_{n}])-\mathbf{K}\mathbb{E}[\mathbf{v}_{n}]=0$, and its covariance is:
(20) $\text{cov}(\mathbf{e}_{n},\mathbf{e}_{n})$ $=\mathbb{E}[((\mathbb{I}-\mathbf{K}\mathbf{C})(\mathbf{A}\mathbf{e}_{n-1}+\mathbf{w}_{n})-\mathbf{K}\mathbf{v}_{n})((\mathbb{I}-\mathbf{K}\mathbf{C})(\mathbf{A}\mathbf{e}_{n-1}+\mathbf{w}_{n})-\mathbf{K}\mathbf{v}_{n})^{T}]$
$=(\mathbb{I}-\mathbf{K}\mathbf{C})(\mathbf{A}\underbrace{\mathbb{E}[\mathbf{e}_{n-1}\mathbf{e}_{n-1}^{T}]}_{\mathclap{\text{cov}(\mathbf{e}_{n-1},\mathbf{e}_{n-1})}}\mathbf{A}^{T}+\underbrace{\mathbb{E}[\mathbf{w}_{n}\mathbf{w}_{n}^{T}]}_{\mathbf{Q}})(\mathbb{I}+\mathbf{K}\mathbf{C})^{T}+\mathbf{K}\underbrace{\mathbb{E}[\mathbf{v}_{n}\mathbf{v}_{n}^{T}]}_{\mathbf{R}}\mathbf{K}^{T}$
where the cross terms are zero because $\mathbf{e}$, $\mathbf{v}$, and $\mathbf{w}$ are assumed to be independent, so $\mathbb{E}[\mathbf{e}\mathbf{w}^{T}]\!=\!\mathbb{E}[\mathbf{v}\mathbf{e}^{T}]\!=\!\mathbb{E}[\mathbf{v}\mathbf{w}^{T}]\!=\!0$. We give this covariance the name $\mathbf{P}$ and note that it is recursively defined like $\mathbf{e}$.
The minimization problem can be further reformulated in terms of this $\mathbf{P}$:
(21) $\hskip 40.00006pt\underset{\mathbf{K}}{\text{minimize}}$ $\text{Tr}[\mathbf{P}_{n}]$
s.t. $\mathbf{P}_{n}=\mathbb{E}[\mathbf{e}_{n}\mathbf{e}_{n}^{T}]$
$\mathbf{e}_{n}=(\mathbb{I}-\mathbf{K}\mathbf{C})\mathbf{A}\mathbf{e}_{n-1}+(\mathbb{I}-\mathbf{K}\mathbf{C})\mathbf{w}_{n}-\mathbf{K}\mathbf{v}_{n}$
because $\mathbb{E}[\|\mathbf{e}\|_{2}^{2}]=\mathbb{E}[\text{Tr}[\mathbf{e}\mathbf{e}^{T}]]$, and the order of trace and expected value can be flipped, since they are linear operators. This form is solvable with calculus [lacey-kf] to determine the optimal $\mathbf{K}$. First manipulate Equation 20:
(22) $\mathbf{P}_{n}$ $=(\mathbb{I}-\mathbf{K}\mathbf{C})\underbrace{(\mathbf{A}\mathbf{P}_{n-1}\mathbf{A}^{T}+\mathbf{Q})}_{\text{name this }\mathbf{P}_{n|n-1}}(\mathbb{I}-\mathbf{K}\mathbf{C})^{T}+\mathbf{K}\mathbf{R}\mathbf{K}^{T}$
$=\mathbf{P}_{n|n-1}-\mathbf{K}\mathbf{C}\mathbf{P}_{n|n-1}-\mathbf{P}_{n|n-1}(\mathbf{K}\mathbf{C})^{T}+\mathbf{K}\mathbf{C}\mathbf{P}_{n|n-1}(\mathbf{K}\mathbf{C})^{T}+\mathbf{K}\mathbf{R}\mathbf{K}^{T}$
$\rightarrow$ $\text{Tr}[\mathbf{P}_{n}]=\text{Tr}[\mathbf{P}_{n|n-1}]\underbrace{-2\text{Tr}[\mathbf{K}\mathbf{C}\mathbf{P}_{n|n-1}]}_{\mathclap{\text{because Tr[$\mathbf{M}$]=Tr[$\mathbf{M}^{T}$] and $\mathbf{P}$=$\mathbf{P}^{T}$}}}+\text{Tr}[\mathbf{K}(\mathbf{C}\mathbf{P}_{n|n-1}\mathbf{C}^{T}+\mathbf{R})\mathbf{K}^{T}]$
where subscript n|n1n|n-1 indicates an a priori estimate at step $n$ based only on information up to step n1n-1, and subscript $n$ (often written $n|n$) indicates the a posteriori estimate after the step resolves and new information is factored in. Using the facts that for matrices $\mathbf{M}$ and $\mathbf{W}$, $(\mathbf{W}\mathbf{M})^{T}=\mathbf{M}^{T}\mathbf{W}^{T}$; $\nabla_{\mathbf{M}}\text{Tr}[\mathbf{M}\mathbf{W}]=\mathbf{W}^{T}$; and $\nabla_{\mathbf{M}}\text{Tr}[\mathbf{M}\mathbf{W}\mathbf{M}^{T}]=\mathbf{M}\mathbf{W}^{T}+\mathbf{M}\mathbf{W}$, which further equals $2\mathbf{M}\mathbf{W}$ when $\mathbf{W}$ is symmetric [duchi-trace-derivative]:
$\rightarrow$ $\nabla_{\mathbf{K}}\text{Tr}[\mathbf{P}_{n}]=-2(\mathbf{C}\mathbf{P}_{n|n-1})^{T}+2\mathbf{K}(\mathbf{C}\mathbf{P}_{n|n-1}\mathbf{C}^{T}+\mathbf{R})=0$
$\rightarrow$ $\mathbf{K}_{\text{opt}}=\mathbf{P}_{n|n-1}\mathbf{C}^{T}(\mathbf{C}\mathbf{P}_{n|n-1}\mathbf{C}^{T}+\mathbf{R})^{-1}\quad\text{\small$\blacksquare$}$
Discrete Kalman filtering, then, is the process of making this closed-form calculation of $\mathbf{K}$ repeatedly at each step and applying it to the system (constraints of Equation 18 or Equation 21) to form updated estimates of the state and error covariance. The framework is flexible enough to accept different system, measurement, and noise covariance matrices at each step, in which case these get subscript $n$, omitted above for brevity. When all matrices are constant, then because the calculation of $\mathbf{K}$ does not rely on the data itself, the filter gain approaches a steady state that accomplishes the same calculation as a Wiener filter [mit-wiener-kalman, wunsch] and can be calculated offline and hard-coded to avoid repeated matrix inversion.
To kick off the filtering process requires: (1) knowledge of the system (what entries of $\mathbf{x},\mathbf{y}$ mean; matrices $\mathbf{A},\mathbf{C}$; and additionally $\mathbf{u},\mathbf{B}$ if there are control inputs or synchronized data streams), (2) knowledge of noise covariances ($\mathbf{Q},\mathbf{R}$), and (3) an initial guess of the state and its covariance ($\mathbf{\hat{x}}_{0},\mathbf{P}_{0}$) to seed the recursion. The filter then proceeds according to the steps of 4, sketched in Figure 21.
Algorithm 4 Kalman Filter
1:assume values for $\mathbf{\hat{x}}_{0},\mathbf{P}_{0}$ (often treated as a priori guesses such that the first two steps of the loop can be skipped on the first iteration)
2:for $n\in\{0,...,N-1\}$ do
3:  calculate predicted next state from past estimate, $\mathbf{\hat{x}}_{n|n-1}=\mathbf{A}\mathbf{\hat{x}}_{n-1}+\mathbf{B}\mathbf{u}_{n}$
4:  propagate the covariance forward a step, $\mathbf{P}_{n|n-1}=\mathbf{A}\mathbf{P}_{n-1}\mathbf{A}^{T}+\mathbf{Q}$
5:  calculate predicted measurement, $\mathbf{\hat{y}}_{n}=\mathbf{C}\mathbf{\hat{x}}_{n|n-1}$
6:  observe true measurement, $\mathbf{y}_{n}$
7:  calculate the current step’s gain, $\mathbf{K}=\mathbf{P}_{n|n-1}\mathbf{C}^{T}(\mathbf{C}\mathbf{P}_{n|n-1}\mathbf{C}^{T}+\mathbf{R})^{-1}$
8:  calculate new state estimate, $\mathbf{\hat{x}}_{n}=\mathbf{\hat{x}}_{n|n-1}+\mathbf{K}(\mathbf{y}_{n}-\mathbf{\hat{y}}_{n})$
9:  calculate new error covariance, $\mathbf{P}_{n}=(\mathbb{I}-\mathbf{K}\mathbf{C})\mathbf{P}_{n|n-1}$ (by simplification of
10:end for                     Equation 22 with $\mathbf{K}_{\text{opt}}$)
[…]
Figure 21: Depiction of the prediction and measurement steps of a Kalman filter. Starting with previous state estimate characterized by $\mathcal{N}(\mathbf{\hat{x}}_{n-1},\mathbf{P}_{n-1})$, first predict the state at the current time using only past information as $\mathcal{N}(\mathbf{\hat{x}}_{n|n-1},\mathbf{P}_{n|n-1})$. The covariance tends to grow over this step due to the addition of $\mathbf{Q}$ in $\mathbf{P}_{n|n-1}$. Then project that prediction into measurement space, $\mathcal{N}(\mathbf{\hat{y}},\mathbf{C}\mathbf{P}_{n|n-1}\mathbf{C}^{T})$. Simultaneously, an unknown true state is corrupted by measurement noise with state-space covariance $\mathbf{\tilde{\Pi}}$ to yield $\tilde{\mathbf{x}}_{n}$ and then projected into measurement space as $\mathcal{N}(\mathbf{y}_{n},\mathbf{R})$. The new state estimate is then formed by optimally—in both the MSE and MAP (section 6.2) senses—combining the prediction with the observation to obtain much tighter probability distributions centered where ellipses overlap. Here both state and measurement space are drawn as 2D, but in general they do not need to have the same dimension.
6.1.1 Linear Modeling Example: Cruise Control
Before delving further into theory, here is a demonstration of the setup required to use a Kalman filter: states, control inputs, measurement outputs, linear evolution, and noise estimates.
Let $\mathbf{y}$ be noisy position data from simulation of a proportional-integral cruise controller that works to maintain a car’s positive constant velocity while driving up and down oscillating hills. To incorporate both the integral feedback control and the action of the hills, state and control inputs are formed as:
$\mathbf{x}$ $=[x_{0},x_{1},x_{2},x_{3}]^{T}=[\text{pos},\text{vel},\text{accel},\text{cumulative pos error}]^{T}$
$\mathbf{u}$ $=[h,v_{d}]^{T}=[\text{hill slope},\text{desired velocity}]^{T}$
and the discrete-time system dynamics are agglomerated into the following matrices:
(23) $\mathbf{A}=\begin{bmatrix}1&\Delta t&\frac{\Delta t^{2}}{2}&0\\ 0&1&\Delta t&0\\ 0&-f_{r}-\frac{k_{p}}{\Delta t}&0&\frac{k_{i}}{\Delta t^{2}}\\ 0&-\Delta t&0&1\\ \end{bmatrix}$, $\mathbf{B}=\begin{bmatrix}0&0\\ 0&0\\ -mg&\frac{k_{p}}{\Delta t}\\ 0&\Delta t\\ \end{bmatrix}$, $\mathbf{C}=\begin{bmatrix}1\\ 0\\ 0\\ 0\\ \end{bmatrix}^{T}$
where $f_{r}$ represents a friction term, $mg$ is the weight of the car, and $k_{p},k_{i}$ are (unitless) proportional and integral feedback gains, respectively. The first two rows of $\mathbf{A}$ essentially integrate position and velocity from acceleration. On the last row of $(\mathbf{A},\mathbf{B})$, estimated and desired velocity at each time step are multiplied by $\Delta t$ to accomplish a simple integration that puts them in units of position, subtracted to create a new position error, and added to the cumulative error. On the third row of $\mathbf{A}$ and $\mathbf{B}$, velocity and position error are divided by $\Delta t$ and $\Delta t^{2}$, respectively, to put them in units of acceleration, then multiplied by appropriate gains. The car’s weight is signed negative so that a positive slope (uphill) will cause deceleration and negative slope (downhill) will cause acceleration. Note that friction due to drag is physically quadratic in velocity, but to use the Kalman filter, we have to assume a linear model, perhaps linearizing around a set point.
A simulation of this system is given in Figure 22, along with Kalman filter and RTS smoothing (section 6.3) estimates using assumed process and measurement noise covariances, $\mathbf{Q}$ and $\mathbf{R}$. Gaussian white noise has been added to data points post hoc, because process and measurement noise contributions at each step combine to produce a single Gaussian, which is the same for all samples if there is constant step size, $\Delta t$. This leaves the underlying noise covariances unspecified and unknown, so we model them according to a common, rough strategy as diagonal matrices of guessed state dimension variances. This assumes independence between state dimensions, which is inaccurate in this case, but the Kalman filter is robust enough to successfully follow the true signal anyway. See section 8.3 for discussion of variable step size and derivation of more sophisticated discrete-time $\mathbf{Q}$.
Figure 22: Simulation of Equation 23 with $\Delta t=0.01$, $mg=10000$, $f_{r}=0.9$, $k_{i}=0.05$, $k_{p}=0.25$, $v_{d}=0.5$, and $h=\big(\sin(2\pi t)+0.3\sin(8\pi t+0.5)+1.2\sin(3.4\pi t+0.5)\big)/100$, without noise (black dotted line) and with additive noise $\sim\mathcal{N}(\mu\!=\!0,\ \sigma\!=\!0.1)$ (blue dots). Using the system matrices along with assumed matrices $\mathbf{P}_{0}=10\cdot\mathbb{I}$, $\mathbf{Q}=1000\cdot\Delta t\cdot\text{diag}[(\frac{1}{2}\Delta t^{2})^{2},\Delta t^{2},1,(\frac{1}{2}\Delta t^{2})^{2}]$, and $\mathbf{R}=[0.1]$, we run a Kalman filter (red line) and RTS smoother (green dashed line) over the noisy points to produce highly accurate estimates of the true signal (root mean squared errors at upper left). The Kalman filter estimate is thrown around by incoming points, especially at the start.
6.2 Bayesian Interpretation of the Kalman Filter
Denoting probability distributions with $p$, the Maximum Likelihood Estimator (MLE) and Maximum A Priori (MAP) estimator find parameters or values, $\boldsymbol{\hat{\theta}}$, given information or measurements, $\mathbf{y}$, according to:
$\boldsymbol{\hat{\theta}}_{\text{MLE}}$ $=\operatorname*{argmax}_{\boldsymbol{\theta}}\ p(\mathbf{y}|\boldsymbol{\theta})$
$\boldsymbol{\hat{\theta}}_{\text{MAP}}$ $=\operatorname*{argmax}_{\boldsymbol{\theta}}\ p(\boldsymbol{\theta}|\mathbf{y})=\operatorname*{argmax}_{\boldsymbol{\theta}}\ p(\mathbf{y}|\boldsymbol{\theta})p(\boldsymbol{\theta})$
where in the latter equation we use Bayes’ rule and the fact its denominator does not depend on $\boldsymbol{\theta}$ and therefore makes no difference to the argmax. The MAP is more powerful because it can account for arbitrary priors, and it maps neatly onto the Kalman context.
The Kalman filter can be viewed as fusing evidence from predictions and sensors. We know information from previous steps as well as current measurements and inputs and now wish to find the best estimate for the present state:
(24) $\mathbf{\hat{x}}_{n}=\operatorname*{argmax}_{\mathbf{x}_{n}}$ $\underbrace{p(\mathbf{x}_{n}|\mathbf{y}_{n},\mathbf{u}_{n},\mathbf{\hat{x}}_{n-1})}$
$=\frac{p(\mathbf{y}_{n}|\mathbf{x}_{n},\mathbf{u}_{n},\mathbf{\hat{x}}_{n-1})p(\mathbf{x}_{n}|\mathbf{u}_{n},\mathbf{\hat{x}}_{n-1})}{p(\mathbf{y}_{n}|\mathbf{u}_{n},\mathbf{\hat{x}}_{n-1})}\quad\text{by Bayes' rule}$
The probability distribution of $\mathbf{y}_{n}=\mathbf{C}\mathbf{x}_{n}+\mathbf{v}_{n},\ \mathbf{v}_{n}\sim\mathcal{N}(0,\mathbf{R})$ is a Gaussian with covariance $\mathbf{R}$ centered at $\mathbf{C}\mathbf{x}_{n}$, which is $\propto e^{\frac{1}{2}\|\mathbf{R}^{-1/2}(\mathbf{y}_{n}-\mathbf{C}\mathbf{x}_{n})\|_{2}^{2}}$. The distribution of $\mathbf{x}_{n}=\mathbf{A}\mathbf{x}_{n-1}+\mathbf{B}\mathbf{u}_{n}+\mathbf{w}_{n},\ \mathbf{w}_{n}\sim\mathcal{N}(0,\mathbf{Q})$ is recursive and cannot be found directly, but we are given $\mathbf{u}_{n}$ and $\mathbf{\hat{x}}_{n-1}$ and thus can construct $\mathbf{\hat{x}}_{n|n-1}=\mathbf{A}\mathbf{\hat{x}}_{n-1}+\mathbf{B}\mathbf{u}_{n}$. Subtracting, $\mathbf{x}_{n}-\mathbf{\hat{x}}_{n|n-1}=\mathbf{A}(\mathbf{x}_{n-1}-\mathbf{\hat{x}}_{n-1})+\mathbf{w}_{n}$, which has known covariance $\mathbf{P}_{n|n-1}$ (derived as first term of Equation 20), so $\mathbf{x}_{n}$ is a Gaussian centered at $\mathbf{\hat{x}}_{n|n-1}$, which is $\propto e^{\frac{1}{2}\|\mathbf{P}_{n|n-1}^{-1/2}(\mathbf{x}_{n}-\mathbf{\hat{x}}_{n|n-1})\|_{2}^{2}}$}. Substituting in Equation 24 and taking the negative logarithm to get rid of exp and turn the maximization into minimization, we obtain:
(25) $\mathbf{\hat{x}}_{n}=\operatorname*{argmin}_{\mathbf{x}_{n}}\|\mathbf{R}^{-1/2}(\mathbf{y}_{n}-\mathbf{C}\mathbf{x}_{n})\|_{2}^{2}+\|\mathbf{P}_{n|n-1}^{-1/2}(\mathbf{x}_{n}-\mathbf{\hat{x}}_{n|n-1})\|_{2}^{2}$
This form is also solvable with calculus. Taking the gradient w.r.t. $\mathbf{x}_{n}$ and setting equal to zero results in $\mathbf{\hat{x}}_{n}=(\mathbf{C}^{T}\mathbf{R}^{-1}\mathbf{C}+\mathbf{P}_{n|n-1}^{-1})^{-1}(\mathbf{C}^{T}\mathbf{R}^{-1}\mathbf{y}_{n}+\mathbf{P}_{n|n-1}^{-1}\mathbf{\hat{x}}_{n|n-1})$, which can be manipulated into $\mathbf{\hat{x}}_{n}=\mathbf{\hat{x}}_{n|n-1}+\mathbf{K}(\mathbf{y}_{n}-\mathbf{C}\mathbf{\hat{x}}_{n|n-1})$, although showing equivalence is tedious,151515unless the Woodbury matrix identity and clever factoring really light your fire see [johns-hopkins].
[…]
(26) $\mathcal{N}(\mathbf{\hat{y}}_{n}\!+\!\mathbf{C}\mathbf{K}(\mathbf{y}_{n}-\mathbf{\hat{y}}_{n}),\ \mathbf{C}\mathbf{P}_{n|n-1}\mathbf{C}^{T}\!-\!\mathbf{C}\mathbf{K}\mathbf{C}\mathbf{P}_{n|n-1}\mathbf{C}^{T})$
$=\mathcal{N}\big(\mathbf{C}(\mathbf{\hat{x}}_{n|n-1}\!+\!\mathbf{K}(\mathbf{y}_{n}-\mathbf{\hat{y}}_{n})),\ \mathbf{C}(\mathbb{I}\!-\!\mathbf{K}\mathbf{C})\mathbf{P}_{n|n-1}\mathbf{C}^{T}\big)$
Projection of random vectors is carried out by a left matrix multiply because $\mathbb{E}[\mathbf{M}\raisebox{0.95pt}{\small$\boldsymbol{\chi}$}]=\mathbf{M}\mathbb{E}[\raisebox{0.95pt}{\small$\boldsymbol{\chi}$}]$, while projection of covariance matrices is done by left and right multiply because $\mathrm{cov}(\mathbf{M}\raisebox{0.95pt}{\small$\boldsymbol{\chi}$},\mathbf{M}\raisebox{0.95pt}{\small$\boldsymbol{\chi}$})=\mathbf{M}\mathrm{cov}(\raisebox{0.95pt}{\small$\boldsymbol{\chi}$},\raisebox{0.95pt}{\small$\boldsymbol{\chi}$})\mathbf{M}^{T}$, so the unprojected $\mathcal{N}(\mathbf{\hat{x}}_{n},\mathbf{P}_{n})$ back in state space can be found by knocking a few $\mathbf{C}$s off Equation 26, leaving exactly the last two update equations of 4.
6.3 Forward-Backward Kalman Smoothing
Traditionally, Kalman filters are designed to run online, in the forward direction, but the state, $\mathbf{\hat{x}}$, can contain abrupt changes due to poor a priori estimates or discordant observations [wunsch]. If the goal is to obtain an accurate derivative offline for an entire data series, then filter estimates can be refined by incorporating information from subsequent steps. Such methods can be thought of as solving the MAP problem (section 6.2) for all steps simultaneously [aravkin, johns-hopkins]:
(27) $\{\mathbf{\hat{x}}_{n}\}=\operatorname*{argmax}_{\{\mathbf{x}_{n}\}}$ $\underbrace{p(\{\mathbf{x}_{n}\}|\{\mathbf{y}_{n}\},\{\mathbf{u}_{n}\})}$
$=\frac{p(\{\mathbf{y}_{n}\}|\{\mathbf{x}_{n}\},\{\mathbf{u}_{n}\})p(\{\mathbf{x}_{n}\}|\{\mathbf{u}_{n}\})}{p(\{\mathbf{y}_{n}\}|\{\mathbf{u}_{n}\})}$
$\propto\Big(\prod_{n=0}^{N-1}\underbrace{p(\mathbf{y}_{n}|\mathbf{x}_{n},\mathbf{u}_{n})}_{=p_{\mathbf{v}}(\mathbf{v}_{n})}\Big)\underbrace{p(\mathbf{x}_{0}|\mathbf{u}_{0})}_{\text{prior}}\prod_{n=1}^{N-1}\underbrace{p(\mathbf{x}_{n}|\mathbf{x}_{n-1},\mathbf{u}_{n})}_{=p_{\mathbf{w}}(\mathbf{w}_{n})}$
where braces represent sets of all states, measurements, or inputs across steps, and $p_{\mathbf{w}},p_{\mathbf{v}}$ are the noise distributions. In this formulation, the prior of each state transition is conditioned on the raw previous state, not the previous state estimate as in the single-step case, because $\mathbf{x}_{n-1}$ is also being optimized over and is therefore available. Considering the joint probability distribution only up to $\mathbf{x}_{n}$ and marginalizing over all $\{\mathbf{x}_{0},...,\mathbf{x}_{n-1}\}$ recovers Equation 24 [johns-hopkins].
The evolution of the system over all steps can be set up as a massive linear equation of stacked vectors and block diagonal matrices. These can be plugged in to create a formulation similar to Equation 25, but using stacked $\mathbf{Q}$ instead of the single-step $\mathbf{P}_{n|n-1}$, because having access to $\mathbf{x}_{n-1}$ enables direct representation of the probability distribution of $\mathbf{x}_{n}$ as a Gaussian with covariance $\mathbf{Q}$ centered at $\mathbf{A}\mathbf{x}_{n-1}+\mathbf{B}\mathbf{u}_{n}$. (See Equation 12 of [aravkin].) Taking the gradient w.r.t. stacked $\{\mathbf{x}_{n}\}$ and setting equal to zero to find the minimum yields a block-tridiagonal linear inverse problem [aravkin, aravkin3, Romberg2016Kalman].
Rauch-Tung-Striebel (RTS) smoothing [RTS, wunsch, aravkin] is a classic algorithm to efficiently solve said system one block at a time, working sequentially backward using information from a forward Kalman filter pass:
(28) $\begin{split}\mathbf{L}_{n}&=\mathbf{P}_{n}\mathbf{A}^{T}\mathbf{P}_{n+1|n}^{-1}\\ \mathbf{\widehat{x}}_{n,\text{RTS}}&=\mathbf{\hat{x}}_{n}+\mathbf{L}_{n}(\widehat{\mathbf{x}}_{n+1,\text{RTS}}-\mathbf{\hat{x}}_{n+1|n})\\ \mathbf{P}_{n,\text{RTS}}&=\mathbf{P}_{n}+\mathbf{L}_{n}(\mathbf{P}_{n+1,\text{RTS}}-\mathbf{P}_{n+1|n})\mathbf{L}_{n}^{T}\end{split}$
Remembering estimates of state and error covariance from every step to support these computations can pose a memory challenge for high-dimensional states or long sequences. When system matrices are constant, $\mathbf{L}$, like $\mathbf{P}$, converges to a steady state, which can optionally be precalculated to save work and memory.
If the system dynamics, $\mathbf{A}$, are invertible, then it is also possible to run the Kalman filter, or whole RTS smoothing, on the data in reverse. Note $\mathbf{Q}$ is left as is for this procedure, because uncertainty due to entropy should still be modeled as increasing across steps, even if time is reversed. Initial guesses $\mathbf{\hat{x}}_{0},\mathbf{P}_{0}$ bias Kalman estimates for several iterations, so combining forward and reverse passes can sometimes improve results at domain edges.
6.4 Generalized Kalman Filtering for Robust Estimation
The Bayesian interpretation of the Kalman filter and the consequent MAP optimization problem (Equations 24 and 27) provide a flexible framework, allowing us to consider alternative noise distributions and distance metrics through modifications of the cost function [aravkin, czech], e.g. via loss functions $V$ and $J$:
(29) $\{\mathbf{\hat{x}}_{n}\}=\operatorname*{argmin}_{\{\mathbf{x}_{n}\}}$ $\sum_{n=0}^{N-1}V(\mathbf{R}^{-1/2}(\mathbf{y}_{n}-\mathbf{C}\mathbf{x}_{n}))$
$+\sum_{n=1}^{N-1}J(\mathbf{Q}^{-1/2}(\mathbf{x}_{n}-\mathbf{A}\mathbf{x}_{n-1}-\mathbf{B}\mathbf{u}_{n}))-\ln(p(\mathbf{x}_{0}))$
For example, assuming measurement noise drawn from a Laplace distribution, $\propto e^{-\|\boldsymbol{\cdot}\|_{1}}$, the corresponding term of the MAP optimization problem (first term of Equations 25 and 29) gets an $\ell_{1}$ norm. Or, if the data has outliers, then because the $\ell_{2}$ norm highly penalizes larger inaccuracies, it can bend the state estimate toward outliers too strongly [ddse]. In this case we may wish to replace the distance metric around the measurement term with something like a Huber loss [Huber1964], which is quadratic only up to some radius, $M$, and then grows linearly:
(30) $\text{Huber}(x,M)=\begin{cases}\frac{1}{2}x^{2}&|x|\leq M\\ M|x|-\frac{1}{2}M^{2}&|x|>M\end{cases}$
Or, in the presence of relatively rare, sudden state changes, the $\ell_{2}$ norm cannot follow fast jumps [ohlsson], so it may be desirable to use the $\ell_{0}$ norm on the process (second term of Equations 25 and 29), or in practice its similarly sparsity-promoting convex relaxation, the $\ell_{1}$ norm.
Figure 23: RTS smoother (green) and robust MAP smoother (dashed pink) estimates on the cruise control data from Figure 22, but this time with 4 randomly chosen points turned into outliers. All matrices remain the same as before except for $\mathbf{Q}$, which has been scaled up by a factor of $10^{5}$ to coerce the solution to deform significantly toward outliers, because RTS smoothing does remarkably well even with outliers when the system matrices are accurate. Despite such a poor choice of $\mathbf{Q}$, a robust smoother with process penalty $\sqrt{2}\|\cdot\|_{1}$ and measurement penalty $1.04\sum\text{Huber}(\cdot,2)$ (constants explained in footnote) is able to ignore outliers and obtain an estimate close to ground truth.
These substitutions can make a big difference to performance, as exhibited in Figure 23, but they typically break the ability to find a closed-form recursive solution by calculus. Fortunately, we can use modern tools like convex optimization to solve the problem efficiently anyway [Boyd_Vandenberghe_2004]—in fact, time complexity is linear thanks to sparse, block-tridiagonal systems [aravkin3, aravkin], similar to the previous section. We can also exploit techniques from this framework to do things like enforce bounds on the state via inequality constraints [aravkin].
There are still limits down this path: Nonlinear models result in nonconvex optimization problems that may or may not be reasonable to solve with alternative approaches, e.g. Gauss-Newton, and some noise distributions, like Student’s t, have merely quasiconvex negative log, necessitating iterative solution procedures [aravkin2, Boyd_Vandenberghe_2004]. Still, this is a powerful generalization when sophisticated system models are available, and we recommend the review given in [aravkin], which includes a tutorial on optimization techniques.
6.4.1 Alternative Notion of Robustness
In controls and signal processing, the term “robust” is also used to mean guaranteed performance under input or model (parametric) uncertainty [robust-control, dullerud2000course, xie]. For unknown inputs, this gives rise to $H_{\infty}$ filtering/smoothing [nagpal, batteries], which is related to $H_{2}$ filtering [robust-control], of which the Kalman filter is the most famous special case. [^17] To cast the Kalman filter into form of Equation 31, the salient output and internal state are both the error $\mathbf{e}$, with dynamics as derived in Equation 19. Vector-valued disturbance can be split into pieces and multiplied by half-powers of covariance matrices $\mathbf{Q}$ and $\mathbf{R}$ to produce appropriately-shaped-but-deterministic (w.r.t. $\mathbf{\tilde{w}}$) process and measurement disturbances: $\begin{bmatrix}\mathbf{e}_{n}\\ \mathbf{e}_{n-1}\end{bmatrix}=\begin{bmatrix}(\mathbb{I}-\mathbf{K}\mathbf{C})\mathbf{A}&\begin{bmatrix}(\mathbb{I}-\mathbf{K}\mathbf{C})\mathbf{Q}^{1/2}&-\mathbf{K}\mathbf{R}^{1/2}\end{bmatrix}\\ \mathbb{I}&0\end{bmatrix}\begin{bmatrix}\mathbf{e}_{n-1}\\ \begin{bmatrix}\mathbf{\tilde{w}}_{\text{proc}}\\ \mathbf{\tilde{w}}_{\text{meas}}\end{bmatrix}_{n}\end{bmatrix}$ Using the time-domain version of the $H_{2}$ norm, which is defined with G’s impulse response (see [robust-control]), along with this ($\mathbf{\tilde{A}},\mathbf{\tilde{B}},\mathbf{\tilde{C}},\mathbf{\tilde{D}}$), we can manipulate the cost function of Equation 32 into $\text{Tr}(\mathbf{\tilde{C}}\mathbf{P}\mathbf{\tilde{C}}^{T})$, where $\mathbf{P}$ is the controllability Gramian $\sum_{m=0}^{\infty}\mathbf{\tilde{A}}^{m}\mathbf{\tilde{B}}\mathbf{\tilde{B}}^{T}(\mathbf{\tilde{A}}^{m})^{T}$, which solves the discrete-time Lyapunov equation, $\mathbf{\tilde{A}}\mathbf{P}_{n-1}\mathbf{\tilde{A}}^{T}-\mathbf{P}_{n}+\mathbf{\tilde{B}}\mathbf{\tilde{B}}^{T}=0$. Algebra recovers Equations 21 and 22, modulo the probabilistic bits. Choosing $\mathbf{\tilde{w}}\sim\mathcal{N}(0,\mathbb{I})$ puts the Gramian sum for $\mathbf{P}$ in the form of an expected value, $\mathbb{E}[\cdot]$, completing the bridge. These methods, perhaps confusingly, are unrelated to MAP generalizations of the Kalman filter. Instead, they consider a linear system named G, which maps exogenous “disturbance” $\mathbf{\tilde{w}}$ and internal state to output error $\mathbf{\tilde{z}}$ and next state (or state derivative in continuous time):
(31) $\begin{bmatrix}\mathbf{\tilde{x}}_{n+1}\\ \mathbf{\tilde{z}}_{n}\end{bmatrix}={\underbrace{\left[\begin{array}[]{c|c}\mathbf{\tilde{A}}&\mathbf{\tilde{B}}\\ \hline\cr\mathbf{\tilde{C}}&\mathbf{\tilde{D}}\end{array}\right]}_{\text{\normalsize G}}}\begin{bmatrix}\mathbf{\tilde{x}}_{n}\\ \mathbf{\tilde{w}}_{n}\end{bmatrix}$
As illustrated in Figure 24, the system can have its own internal logic, such as a “plant” with self-contained state and linear dynamics connected to a controller (or observer). Signals in this framework are typically taken to be $\in L_{2}$ so they have nice inner products, and the complex-valued transfer function that maps between signals is constrained to a Hardy space, $H_{p}$, to ensure its norm integrals are bounded.
[…]
Figure 24: In the framework of H norms, system G takes in a disturbance signal and produces an error signal. Internally, G can have constituent components that affect its behavior, which are factored into the overall Equation 31.
We then form the following optimization problem:
(32) $\operatorname*{argmin}_{\mathbf{K}}$ $\|\text{G}\|_{H_{p}}$
s.t. G’s dynamics satisfied
where $\mathbf{K}$ are tunable parameters in the system and $\|\cdot\|_{H_{p}}$ denotes a Hardy norm. We care about $p=2$, because $\|\text{G}\|_{H_{2}}$ is the inner product of G’s transfer function with itself in Hardy space, which integrates the response to sinusoidal inputs over all frequencies, thereby penalizing the average response [robust-control]. We also care about $p=\infty$, where $\|\text{G}\|_{H_{\infty}}\triangleq\sup_{\omega}\sigma_{\text{max}}(\text{G})$, [^18] In practice, this cannot be solved directly, but the Kalman–Yakubovich–Popov lemma provides a way to check whether $\|\text{G}\|_{H_{\infty}}<\gamma$ using a Linear Matrix Inequality [boyd-lmis], which is a convex problem. If infeasible, raise $\gamma$ to loosen, and if feasible, step $\gamma$ down to tighten; it’s quasiconvex. which penalizes the worst-case response, the system’s most sensitive mode at its most sensitive frequency [robust-control]. A result from functional analysis says the $H_{\infty}$ norm also gives $\sup_{\mathbf{\tilde{w}}\neq 0}\frac{\|\mathbf{\tilde{z}}\|_{2}}{\|\mathbf{\tilde{w}}\|_{2}}$, i.e. the maximal ratio of output magnitude (error) to input magnitude (disturbance), including arbitrary, non-sinusoidal disturbances [robust-control], making it a useful design criterion when error amplification must never exceed a hard bound.
However, this mathematically intense framework is often weaker than desired: $H_{p}$-optimal solutions are optimal regardless of any particular form of bounded-energy disturbance. This can be counterintuitive, especially because it treats disturbances as deterministic, whereas from the outset we have considered random noise (section 5). The classic Kalman filter is always $H_{2}$-optimal because it minimizes the system norm, but we understand it is also the “best” estimator in a probabilistic sense under certain conditions and have seen examples in section 6.4 of extending that same notion of best (MAP) to alternative assumptions, requiring filter modifications. Thus this $H_{2}$ sense of optimality may not actually be best for a given scenario. Similarly, the $H_{\infty}$ filter is impressively robust but extremely conservative. Some systems that can be treated with Kalman filtering have unbounded $H_{\infty}$ norm, making $H_{\infty}$ filter design infeasible. Even when applicable, an $H_{\infty}$ filter places more weight on measurements and less on priors than a Kalman filter, allowing more high-frequency noise through [crassidis_junkins]. This notion of robustness is not the right tool for model-based filtering unless truly nothing is known about the disturbance, a rare situation. A smoother estimate, more appropriate for differentiation, will be possible under very bare assumptions, like wide uniform priors and anticipation of outliers.
6.5 Nonlinear Filtering
The methods covered so far are all limited to linear models, but many real-world systems evolve according to nonlinear equations. This can even include linear systems with nonlinear parametric dependency if parameters are estimated as part of the state. For example, say the gain terms $k_{p}$ and $k_{i}$ in the cruise control example from section 6.1.1 are unknown, then the system can be considered to have the following state and dynamics:
$\mathbf{x}=[x_{0},x_{1},x_{2},x_{3},k_{i},k_{p}]^{T}$
$\mathbf{x}_{n}$ $=\boldsymbol{f}(\mathbf{x}_{n-1},\mathbf{u}_{n-1})=\begin{bmatrix}x_{0}+\Delta t\ x_{1}+\frac{\Delta t^{2}}{2}x_{2}\\ x_{1}+\Delta t\ x_{2}\\ -f_{r}x_{1}-mg\ u_{0}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{k_{i}}{\Delta t^{2}}x_{2}}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{k_{p}}{\Delta t}u_{1}}\\ x_{4}+\Delta t\ u_{1}\\ k_{i}\\ k_{p}\end{bmatrix}_{n-1}$
$\mathbf{y}_{n}$ $=\boldsymbol{h}(\mathbf{x}_{n-1},\mathbf{u}_{n-1})=[x_{1}]_{n-1}$
where blue font indicates the nonlinear terms.
Addressing nonlinear systems in the Kalman framework requires mathematical extension. The oldest and simplest approach is fittingly called the Extended Kalman Filter (EKF), which forward-solves the process dynamics given by $\boldsymbol{f}(\mathbf{x},\mathbf{u})$ to predict the a priori state estimate and uses the measurement function, $\boldsymbol{h}(\mathbf{x},\mathbf{u})$, to get the current measurement, but then uses Jacobians [^19] If $\boldsymbol{f}$ and $\boldsymbol{h}$ are linear functions, expressible as products of coefficient matrices and variable vectors, then their Jacobians are simply the matrices, and the EKF reduces to the KF. evaluated at the current state estimate in place of system matrices in the Kalman filter equations to find error covariance, $\mathbf{P}$, and gain, $\mathbf{K}$:
$\mathbf{A}_{n}=\frac{\partial\boldsymbol{f}(\mathbf{x},\mathbf{u})}{\partial\mathbf{x}}\Bigg|_{\hat{\mathbf{x}}_{n-1},\mathbf{u}_{n}},\quad\mathbf{B}_{n}=\frac{\partial\boldsymbol{f}(\mathbf{x},\mathbf{u})}{\partial\mathbf{u}}\Bigg|_{\hat{\mathbf{x}}_{n-1},\mathbf{u}_{n}},\quad\mathbf{C}_{n}=\frac{\partial\boldsymbol{h}(\mathbf{x},\mathbf{u})}{\partial\mathbf{x}}\Bigg|_{\hat{\mathbf{x}}_{n|n-1},\mathbf{u}_{n}}$
Unlike the linear Kalman filter, the EKF is not optimal and not guaranteed to converge; its first-order linearizations may not be able to keep up with highly nonlinear dynamics.
Further innovations have been made to better handle acute nonlinearities, the most famous being the Unscented Kalman Filter (UKF) [julier-uhlmann, julier-uhlmann2], which predicts subsequent error covariance [^20] In the original UKF, the error covariance matrix, $\mathbf{P}$, can lose positive definiteness due to numerical issues. This can be largely resolved by using the square-root implementation, which guarantees the error covariance remains positive definite by estimating and updating the Cholesky factor, $\mathbf{S}$ such that $\mathbf{P}=\mathbf{SS}^{T}$, instead of $\mathbf{P}$ itself [merwe2001sqrt]. and state by propagating specially-chosen (“sigma”) points through the actual nonlinear dynamics. This can be thought of as a kind of particle filter [russell-norvig], but the UKF enforces a Gaussian state estimate and exploits this structure to make next estimates using a linear (in the state dimension) number of points, whereas a generic particle filter requires an exponential number of particles to represent arbitrary probability distributions, per the curse of dimensionality. Figure 25 offers a visual comparison. This method “captures the posterior mean and covariance accurately to the $3^{\text{rd}}$ order for any nonlinearity.” [oregon]. For improved estimates in the post hoc regime, an RTS-like smoothing algorithm based on the UKF was derived by [sarkka].
Figure 25: From [oregon]. Left: Particles sampled from a Gaussian are passed through a nonlinearity to produce a non-Gaussian distribution with calculable mean and covariance. Middle: The EKF uses a local linear approximation of the nonlinearity to propagate mean and covariance inaccurately. Right: Specially-chosen points are passed through the dynamics, and the results are weighted according to the UKF scheme to produce a much better approximation of the true mean and covariance, although non-Gaussianity of the final distribution is not captured.
6.6 Recommendations
Due to Gauss-Markov theorem, the standard Kalman filter is the Best Linear Unbiased Estimator (BLUE) for noise distributions with zero mean and known covariance that are uncorrelated through time (white) and independent of the system state (homoscedastic) [aravkin, Romberg2016Kalman]. For linear systems, the Kalman filter is also the most efficient unbiased estimator of any kind, as the error covariance converges to the Cramér-Rao lower bound [crassidis_junkins, Cramer1946]. Coupling these facts with the central limit theorem, which states the combined influence of many random sources tends toward a Gaussian, the standard Kalman filter is surprisingly effective across a wide array of practical contexts.
Even so, the classical approach has limits: If data is contaminated with outliers or non-Gaussian noise, or when the system does not evolve linearly, the most optimal estimator becomes biased or nonlinear. Sophisticated generalizations and extensions of Kalman filtering can be complicated to set up and solve, especially as concerns like nonlinearity and robustness stack together, making the mathematics ever more involved. This remains an active area of research with a dizzying number of filter variants [chilean, friston] and novel approaches [martin, variational-kalman].
At an even more basic level, system modeling can be a challenge, and a model-based approach may not confer an advantage versus the model-free paradigm (section 7), depending on accuracy of the model. Automatic model discovery techniques exist but are difficult to apply to the differentiation problem, because they themselves rely on good derivative estimates [sindy, ddse]. In some cases, blending a partially-known model with the performance metrics of section 7.1 may help in estimating unknown parameters like noise levels.
7 Differentiating Noisy Data Without Prior Knowledge
In the absence of a model, estimating the derivative of a function, x(t)x(t), [^21] It is common in the context of noisy data to keep with the Kalman filter notation of section 6, where the symbol $y$ is reserved for noisy measurements and underlying signal is called $x$, so the independent variable must take on another symbol, $t$. This contrasts with earlier sections of this review, where functions are usually named $y$, taking input $x$. given noisy samples, $\mathbf{y}=\mathbf{x}+\boldsymbol{\eta}=x(\{t_{n}\})+\eta(\{t_{n}\})$, where $\{t_{n}\}$ is a set of sample points and $\eta$ is a noise process, is an ill-posed problem. In other words, the data are ambiguous, so it is not possible to solve for the derivative without imposing prior assumptions. Nevertheless, practitioners must solve this ill-posed problem routinely, often without much guidance. This has been increasingly true in data-driven science, where large-scale and often high-dimensional data streams are used to construct models [ddse, apriori-boulder].
7.1 Performance Metrics in the Presence and Absence of Ground Truth
A whole cornucopia of methods have been developed to estimate derivatives of noisy data, each with a set of tunable hyperparameters that affect its behavior, such as polynomial degree, window width, and number of iterations, which we collectively call $\Phi$. Differentiating in the absence of prior knowledge comes down to choosing a method and hyperparameter settings that achieve a good answer, where “good” entails two major concerns: closeness of fit and complexity [apriori-boulder]. Following [floris2020], we use the following metrics:
  1. 1.
    Accuracy: The derivative estimate, $\mathbf{\hat{\dot{x}}}$, should be faithful to the true derivative, $\mathbf{\dot{x}}$, resulting in a small root mean squared error:
    (33) $\text{RMSE}=\sqrt{\frac{1}{N}\sum{(\mathbf{\hat{\dot{x}}}-\mathbf{\dot{x}})^{2}}}$
  2. 2.
    Bias: The derivative estimate should be unbiased in the sense its accuracy should not depend on the underlying true value of the derivative. Bias occurs in low-pass filtering, for instance, because it not only removes noise but also dulls sharp slopes, causing systematic under-prediction of larger-magnitude derivative values. We can measure bias with the correlation coefficient between estimate error, $\mathbf{\hat{\dot{x}}}-\mathbf{\dot{x}}$, and the actual derivative:
    (34) $R^{2}=\frac{\text{cov}(\mathbf{\hat{\dot{x}}}-\mathbf{\dot{x}},\mathbf{\dot{x}})}{\text{var}(\mathbf{\hat{\dot{x}}}-\mathbf{\dot{x}})\text{var}(\mathbf{\dot{x}})}$
Both the above metrics require knowledge of the true derivative, $\mathbf{\dot{x}}$, which in real-world applications is usually unknown, so the authors of [floris2020] suggest a proxy loss function that uses only measurements $\mathbf{y}$ and the assumption that noise is largely band-separable from signal (section 5) to balance fidelity and smoothness:
(35) $L(\Phi)=\text{RMSE}\big(\text{trapz}(\mathbf{\hat{\dot{x}}}(\Phi))+\mu,\ \mathbf{y}\big)+\gamma\text{TV}\big(\mathbf{\hat{\dot{x}}}(\Phi)\big)$
where vector $\mathbf{\hat{\dot{x}}}(\Phi)$ is the estimate returned by a differentiation method parameterized by $\Phi$; trapz is numerical integration with trapezoidal rule, which is less sensitive to noise than higher-order quadrature rules [pynumdiff]; $\mathbf{y}$ are observed noisy measurements; $\mu$ corrects for a lost constant of integration by solving $\operatorname*{argmin}_{\mu}\|\text{trapz}(\mathbf{\hat{\dot{x}}})+\mu-\mathbf{y}\|_{2}$, which is just the mean difference between the two vectors; TV is normalized Total Variation (section 7.6), equal to $\frac{1}{N}\|\mathbf{\hat{\dot{x}}}_{0:N-1}-\mathbf{\hat{\dot{x}}}_{1:N}\|_{1}$; and $\gamma$ is a hyperparameter, with larger values of $\gamma$ favoring smoother derivative estimates.
In the presence of outliers, RMSE can weight spurious points too highly, so we further suggest a generalization of Equation 35 using the Huber loss (Equation 30):
(36) $L(\Phi)=\sqrt{\frac{2}{N}\sum\text{Huber}\Big(\text{trapz}(\mathbf{\hat{\dot{x}}}(\Phi))+c-\ \mathbf{y},M\sigma_{\text{MAD}}\Big)}+\gamma\text{TV}\big(\mathbf{\hat{\dot{x}}}(\Phi)\big)$
where Huber is applied pointwise to vector inputs; $\sigma_{\text{MAD}}$ is a measure of scatter in the residuals (first Huber argument) using median absolute deviation (MAD), normalized by the median absolute value of the standard normal distribution, $\mathcal{N}(0,1)$, so its scale matches that of standard deviation for Gaussian errors; $M$ parametrizes where the Huber loss changes from quadratic to linear, in units of $\sigma_{\text{MAD}}$; and $c$ is an integration constant, this time estimated robustly by solving $\operatorname*{argmin}_{c}\sum\text{Huber}(\text{trapz}(\mathbf{\hat{\dot{x}}})+c-\mathbf{y},M\sigma_{\text{MAD}})$. As $M\to\infty$, Equation 36 reduces to Equation 35, but for even moderate $M$ the vast majority of data points are treated with RMSE; e.g., assuming Gaussian inliers, the portion beyond $M\sigma$ can be found with 2*(1-scipy.stats.norm.cdf(M)), which is about 0.0455 for $M=2$ and 0.0027 for $M=3$.
Due to the complex relationship between $\Phi$ and $\mathbf{\hat{\dot{x}}}$ through a differentiation method, Equations 35 and 36 are unfortunately not convex in $\Phi$, nor are gradients with respect to $\Phi$ easy to find. But they can still be optimized via strategies such as Nelder-Mead [nelder-mead], a downhill simplex direct search, starting from multiple initial conditions [pynumdiff, floris2020].
[…]
Choosing $\gamma$ to achieve both low RMSE and $R^{2}$ is found to depend on the temporal resolution and frequency content of the data and not on function shape, noise level, or differentiation method, and a heuristic formula is provided to determine its value [floris2020]:
(37) $\ln(\gamma)=-1.6\ln(f)-0.71\ln(\Delta t)-5.1$
where $f$ is the signal bandlimit in Hz, i.e. the highest frequency expected to come from signal rather than from noise, and the constants come from a best-fit regression. The idea is that when a signal contains higher frequencies, its derivative cannot be smoothed as aggressively.
Armed with the knowledge that smoothing is indeed the right approach, the methods in this section can broadly be thought of as implementing smoothing in a number of different ways.
7.2 Prefiltering Methods
The simplest way to smooth data is as an independent first step, before calculating a derivative. This is a heavy-handed operation, so although global differentiation methods offer an error-bound advantage versus local ones [ahnert], prefiltering erases a lot of information and nullifies the difference. As such, simple Finite Difference is ideal for this case, cheap to compute and imposing no extra assumptions.
A common method of smoothing is to simply take a moving average, which is equivalent to convolving with a discrete kernel of uniform values that sum to 1. We can also take the median over windows, which can remove spurious values while preserving sharp transitions if noise levels are not too high, or we can convolve with alternative normalized kernels sampled from Gaussian or Friedrichs222222a bump function, e.g. $e^{-1/(1-x^{2})}$, $x\in(-1,1)$ functions. These kernel methods constitute crude digital low-pass filters, with frequency responses (magnitude Bode plots [oppenheim]) shown in the first three panels of Figure 27. The median filter is not visualized, because it is nonlinear and lacks a well-defined response.
Figure 27: Frequency responses of possible prefilters: moving average (top left), Gaussian kernel (top right), Friedrichs kernel (lower left), and Butterworth filter (lower right). Plots are generated by scipy.signal.freqz, with frequency given as a fraction of the Nyquist rate by scaling angular frequency $\omega\in[0,\pi]$ by $\frac{1}{\pi}$. Frequency caps at the Nyquist rate, $\frac{f_{s}}{2}$, because higher frequencies alias down into this band (as in Figure 18). All methods demonstrate low-pass character, but only Butterworth is purpose-designed, while the others exhibit undesirable humps in higher frequencies. The bandwidth is said to stop where power output is halved, about -3 decibels per $10\log_{10}(\frac{1}{2})$.
Figure 28: Derivative estimates for noisy cruise control data (Figure 22) found by applying a kernel in a sliding window followed by second-order Finite Difference, using three different choices of hyperparameters.
Instead of relying on such approximations, we can design and apply a true low-pass filter directly. A simple option is the Butterworth filter, with gain (complex transfer function magnitude) given by:
(38) $G(\omega)=|H(i\omega)|=\frac{1}{\sqrt{1+\big(\frac{\omega}{\omega_{c}}\big)^{2n}}}$
where $\omega$ is frequency, $\omega_{c}$ is the “cutoff” where power is halved, and $n$ is the filter order. This response, visualized in Figure 27, is intended to stay as flat as possible in the passband, thereby introducing minimal distortion. In contrast to the global Fourier-Spectral Method (section 4.2.1), which can be used as an ideal low-pass filter by zeroing out higher modes followed by inverse transforming, the Butterworth filter is local and can only achieve gradual cutoff. But as compensation, Butterworth filters are applicable to any signal, not only those that are periodic.
Figure 29: Derivative estimates for noisy cruise control data (Figure 22) found by Butterworth smoothing with three different choices of hyperparameters followed by second-order Finite Difference.
7.3 Iterated Finite Difference
In the presence of noise, Finite Difference’s division by powers of the (small) step size, as in the formulas of Table 2, greatly amplifies deviations in noisy $\mathbf{y}$, resulting in poor derivative estimates of the underlying signal, $\mathbf{x}$. In addition, forward- and backward-differencing equations tend to have large constants, which magnify noise and cause characteristic blowup at the domain edges, as seen in Figure 30.
Figure 30: Second-Order Finite Difference on the cruise control data of section 6.1.1. The unmodified method has no mechanism to estimate smoothed $\mathbf{\hat{x}}$ from the data, resulting in exceptionally jagged derivative estimates. Domain edges are found with formulas that further intensify noise.
The first problem, division by $\Delta t$, can be counterbalanced by performing integration, which multiplies by $\Delta t$. These operations may appear to pointlessly cancel out; indeed the most naive scheme, first-order Finite Difference followed by cumulative summation, essentially just shifts the data:
$\mathbf{y}=[y_{0},y_{1},y_{2},y_{3}]\rightarrow\text{first-order-FD}\big(\mathbf{y}\big)$ $=[y_{1}-y_{0},y_{2}-y_{1},y_{3}-y_{2},y_{3}-y_{2}]/\Delta t$
$\rightarrow\text{cumsum}\big(\text{first-order-FD}\big(\mathbf{y}\big)\big)\cdot\Delta t$ $=[y_{1},y_{2},y_{3},y_{3}]-y_{0}$
However, second-order center-difference followed by cumulative trapezoidal integration (trapz) introduces a factor of $\frac{1}{2}$. Ignoring the edges, we find:
$\text{second-order-FD}\big([y_{0},y_{1},y_{2},y_{3}]\big)[n=1,2]$ $=[y_{2}-y_{0},\ y_{3}-y_{1}]/2\Delta t$
[…]
In the language of digital filters, this is effectively implementing the infinite impulse response (IIR) filter, $\hat{x}[n]=\hat{x}[n-1]+\frac{1}{4}(y[n+1]+y[n]-y[n-1]-y[n-2])$, where, in a stunning reversal of ordinary digital filtering notation, $y$ are inputs and $\hat{x}$ are outputs. This is yet another low-pass filter, with response shown in Figure 31. Iterating the procedure is like passing through this filter multiple times.
Figure 31: The frequency response of the IIR filter equivalent to one round of second-order center-differencing followed by trapezoidal integration. Iterating more times causes the curve to multiply itself and bend down more rapidly, reaching -80dB by about 1/3 Nyquist for 30 repetitions. The passband is also curved, not flat.
Higher-order differencing schemes smooth across more samples, although the center-differencing coefficients of immediately-neighboring samples tend to become greater, concentrating energy more locally and necessitating more iterations to adequately smooth. It is also possible to use higher-order quadrature rules, but even Simpson’s 1/3 rule, which integrates over $2\Delta t$ at a time with $\frac{1}{3}(y_{n-2}+4y_{n-1}+y_{n})$, contains a coefficient >1>\!\!1, which risks amplifying noise.
Edge blowup can be avoided by simply using lower-order differencing equations, but one can also solve the Finite Difference linear inverse problem (Equation 2) with a spaced-out stencil, e.g. $[0,2,4]$, to yield coefficients that do not exceed unity.
Figure 32: Iterated Finite Difference on noisy cruise control data (Figure 22) with three different choices of hyperparameters.
7.4 Polynomial Fits
This section covers methods that set up and solve regression problems to approximate data with smooth polynomial functions, which can then be differentiated analytically by power rule and evaluated at sample points.
7.4.1 Sliding Window Polynomial Fits
A function $x(t_{i})=c_{0}+c_{1}t_{i}+...+c_{d}t_{i}^{d}$, where $d$ is some degree, can be fit to data $\mathbf{y}$ by solving the linear inverse problem:
(39) $\begin{bmatrix}1&t_{1}&t_{1}^{2}&\cdots&t_{1}^{d}\\ 1&t_{2}&t_{2}^{2}&\cdots&t_{2}^{d}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 1&t_{n}&t_{n}^{2}&\cdots&t_{n}^{d}\end{bmatrix}\begin{bmatrix}c_{0}\\ c_{1}\\ \vdots\\ c_{d}\end{bmatrix}\approx\mathbf{y}$
If the Vandermonde matrix is square, the two sides of Equation 39 can be made equal, and $x(t)$ goes through all data points. To obtain a smoother function, we limit degree $d$, making the system overdetermined, and minimize the $\ell_{2}$ norm of the difference between the two sides.
However, neither a low nor high choice of $d$ may be acceptable, because low-degree polynomials have very little expressive power while higher-degree fits tend to exhibit Runge’s phenomenon (Figure 9). To circumvent this, we can restrict a fit’s scope to some local subset of data and keep degree fairly low, because local subsets do not have as much opportunity for complex behavior. Covering the whole domain then requires many such fits, which can be thought of as happening inside a sliding window that moves by some stride, as illustrated in Figure 33.
Figure 33: Polynomial fits, each with its own color, to batches of 40 samples with a stride of 20, on cruise control data from section 6.1.1.
Samplings of these fits and their derivatives can then be combined by averaging their overlaps or, if a kernel is applied to weight fit samples, weighted averaging. Example outputs are shown in Figure 34
Figure 34: Sliding-window polynomial differentiation of noisy cruise control data (Figure 22) with three different choices of hyperparameters.
7.4.2 Savitzky-Golay Filtering
In the special case samples are equispaced, no weighting kernel is desired, and each output is found as a single sample from its own fit (moving-polynomial fit, stride of 1), output values can be calculated without explicitly solving linear inverse problems, because the solution can be written as a fixed linear combination of window inputs [savitzky1964smoothing, schafer-savgol].
This is seen by an argument very similar to the derivation of Finite Difference equations in section 4.1. Specifically, the moving window implies a stencil of $S$ samples, $s_{0},s_{1},...,s_{S-1}]$, e.g. [1,0,1][-1,0,1] to sample 3 equispaced points centered on a point of interest. We then wish to find a $d^{\text{th}}$ order polynomial of the form $c_{0}+c_{1}s+c_{2}s^{2}+...+c_{d}s^{d}$ that when sampled at the stencil-points approximates the data, $\mathbf{y}_{\text{window}}$, at the corresponding points. Paralleling Equations 2 and 39, this can be set up as:
$\begin{bmatrix}s_{0}^{0}&\cdots&s_{0}^{d}\\ \vdots&\ddots&\vdots\\ s_{S-1}^{0}&\cdots&s_{S-1}^{d}\end{bmatrix}\begin{bmatrix}c_{0}\\ \vdots\\ c_{d}\end{bmatrix}=\begin{bmatrix}|\\ \mathbf{y}_{\text{window}}\\ |\end{bmatrix}$
At the point of interest, $s=0$, the fit’s value is given by $c_{0}$ alone, which can be found as the dot product of $\mathbf{y}_{\text{window}}$ with the first row of the pseudoinverse of the stencil Vandermonde matrix. Because said matrix depends only on the chosen stencil and polynomial degree $d$, the coefficients multiplying $\mathbf{y}_{\text{window}}$ are independent of the data and can be found offline. Moreover, the vector of coefficients constitutes a digital filter, with frequency response shown in Figure 35, which can be applied to data efficiently via convolution. [^23] Data edges, where the stencil may not fit, must be treated specially, often by extending the data with repeats or mirroring prior to convolution or by explicitly calculating edge polynomial fits so they can be evaluated beyond single points.
Figure 35: The frequency response of a Savitzky-Golay filter. It is very flat in the passband, which minimizes distortion [schafer-savgol]. When polynomial degree is 1, it is identical to the moving average filter (Figure 27).
Savitzky-Golay can also deliver values of the derivative, by
$\frac{d}{dt}(c_{0}+c_{1}s+...+c_{d}s^{d})=\frac{d}{ds}(c_{0}+c_{1}s+...+c_{d}s^{d})\frac{ds}{dt}=\frac{c_{1}}{\Delta t}$
because $s=\frac{t}{\Delta t}$ is just a scaled variable, so $\frac{ds}{dt}=\frac{1}{\Delta t}$; and $c_{1}$, like $c_{0}$, is also merely a dot product of $\mathbf{y}_{\text{window}}$ with a vector of coefficients that can be found offline.
In practice, this method’s results can be somewhat thrown around by the independence of implicit fits at subsequent data points, so it can be helpful to smooth the results in a postprocessing step by convolving with a Gaussian kernel.
Figure 36: Savitzky-Golay differentiation followed by Gaussian smoothing of noisy cruise control data (Figure 22) with three different choices of hyperparameters.
7.4.3 Spline Smoothing
Splines are piecewise polynomials, constructed to be continuous and to have continuous derivatives up to some order. Joins between the pieces are known as knots. With a sufficient number of knots, a spline can achieve a high degree of expressive power, but because each piecewise component only needs to consider local information, it can be fit with a low-degree polynomial, which averts Runge’s phenomenon (Figure 9).
Each component of a spline can be given a definite formula, for example a cubic:
$S_{k}(t)=c_{k,0}+c_{k,1}(t-t_{k})+c_{k,2}(t-t_{k})^{2}+c_{k,3}(t-t_{k})^{3},\quad t\in[t_{k},t_{k+1}]$
By enforcing continuity of the function and its first derivative at the left and right boundary points of each spline component, we get the following four equations to determine the four unknown coefficients:
$\left\{\begin{aligned} S_{k}(t_{k})&=S_{k-1}(t_{k})\\ S_{k}(t_{k+1})&=S_{k+1}(t_{k+1})\\ \dot{S}_{k}(t_{k})&=\dot{S}_{k-1}(t_{k})\\ \dot{S}_{k}(t_{k+1})&=\dot{S}_{k+1}(t_{k+1})\end{aligned}\right$.
At the left and right boundaries of the domain, a neighboring component is no longer available in one direction, so alternative constraints must be imposed, such as $\ddot{S}_{0}(t_{0})=0$ and $\ddot{S}_{n-1}(t_{n})=0$. All these equations can ultimately be written as a sparse linear inverse problem with target $\mathbf{y}$, similar to Equation 5. If there are as many knots as sample points, the spline intersects every data point, which can be useful for refining coarsely-spaced, noiseless data, as shown in Figure 37(a).
[…]
Figure 37: (a) Cubic spline fitting to noiseless data for refinement of a grid from $\Delta t$ to $\Delta\tau$. (b) Cubic spline smoothing of noisy data. In this case, the spline is not forced to go through the samples, rather a balance between data fitting and smoothness is sought.
For noisy data, the spline-fitting problem can be modified so the solution passes between data points, as depicted in Figure 37(b). There are a couple different ways to do this. The first [wahba1975smoothing, craven1978smoothing] is to add a term that penalizes the integral of the second derivative:
(40) $\operatorname*{argmin}_{\boldsymbol{\theta}}\sum_{n=0}^{N-1}(y_{n}-S(t_{n};\boldsymbol{\theta}))^{2}+\lambda\!\!\int\limits_{t_{0}}^{t_{N-1}}\!\!\!S^{\prime\prime}(u;\boldsymbol{\theta})du$
where $S$ is the overall spline, $\boldsymbol{\theta}$ represents all coefficients collected together, and $\lambda$ is a hyperparameter which balances fitting the data with keeping the second derivative small. The second possibility [dierckx1975] is to simply require the sum of squares of the best fit be within a bound, $s$, which corresponds to the minimization of Equation 40 with $\lambda=0$ and an additional condition:
(41) $\text{s.t.}\quad\sum_{n=0}^{N-1}(y_{n}-S(t_{n}))^{2}\leq s$
where larger $s$ allows for more wiggle room and therefore more smoothness. This will be either feasible or infeasible, depending on the number of knots.
One route to solve these involves representing $S$ as a sum of cubic B-splines, B for “basis”.
(42) $S(t)=\sum_{j=1}^{m}\alpha_{j}B_{j}(t)$
where $B_{j}(t)$ is the $j^{\text{th}}$ B-spline, $\alpha_{j}$ is a coefficient, and $m$ is the total number of basis functions, which depends on the number of knots and smoothness we wish to achieve. B-splines, visualized in Figure 38, are designed to be nonzero over short intervals, redolent of the basis functions from the Finite Element Method (footnote 5).
Figure 38: Visualization of cubic basis splines. Knot locations are shown as vertical dashed lines. Each basis function has compact support, i.e. takes nonzero value only on a number of between-knot intervals at most equal to the spline’s degree. Within each between-knot interval, the basis splines form a partition of unity, i.e. sum to one.
Equation 42 can be substituted in Equation 40 to give:
$\operatorname*{argmin}_{\boldsymbol{\alpha}}\sum_{n=0}^{N-1}\Big(y_{n}-\sum_{j}^{m}\alpha_{j}B_{j}(t_{n})\Big)^{2}+\lambda\!\!\int\limits_{t_{0}}^{t_{N-1}}\!\!\!\Big(\sum_{j}^{m}\alpha_{j}B_{j}^{\prime\prime}(u)\Big)^{2}du$
where $t_{n}$ is a sample point, not a knot point as sometimes denoted when working with splines. This can be further written in matrix form:
(43) $\operatorname*{argmin}_{\boldsymbol{\alpha}}\|\mathbf{y}-\mathbf{B}\boldsymbol{\alpha}\|^{2}+\lambda\,\boldsymbol{\alpha}^{T}\mathbf{R}\boldsymbol{\alpha}$
where $\mathbf{B}\in\mathbb{R}^{N\times m}$ has entries $B_{nj}=B_{j}(t_{n})$ and $\mathbf{R}\in\mathbb{R}^{m\times m}$ is the roughness penalty matrix, with entries $R_{jk}=\int B_{j}^{\prime\prime}(\tau)B_{k}^{\prime\prime}(\tau)\ d\tau$.
Taking the gradient with respect to $\boldsymbol{\alpha}$ and setting equal to zero yields a linear inverse problem:
$(\mathbf{B}^{T}\mathbf{B}+\lambda\mathbf{R})\boldsymbol{\alpha}=\mathbf{B}^{T}\mathbf{y}$
Due to the compactness of B-splines, only $O(d)$ can overlap any point in the domain, where $d$ is the degree of $S$, and therefore each row of $\mathbf{B}$ contains only $O(d)$ nonzero elements, so $B^{T}B$ can be found in $O(Nd^{2})$ instead of $O(Nm^{2})$. Both $\mathbf{B}^{T}\mathbf{B}$ and $\mathbf{R}$ are banded, i.e. nonzero on only $O(d)$ diagonals, allowing the system to be solved efficiently in $O(md^{2})$.
If using smoothness hyperparameter $s$ instead of $\lambda$ (Equation 41), the second term of Equation 43 disappears, breaking regularization. However, smoothness can still be imposed by keeping the number of knots low, because a spline with fewer knots has less expressive power. So we can find a good choice of parameters by refining a global, low-degree polynomial fit by iteratively adding knots at locations of greatest disagreement between the spline and data. [^24] This is the approach favored by scipy [scipy-smoothing-splines].
In practice, it can be beneficial to repeat the spline-fitting process a few times, replacing the data with spline samples after each iteration, to gradually remove noise while preserving data peaks, because splines can tend to oversmooth if forced to use too few knots right upfront.
Figure 39: Spline differentiation of noisy cruise control data (Figure 22) with three different choices of hyperparameters.
7.5 Basis Fits
This section covers methods that build function approximations from smooth basis functions, both global and local. B-splines, presented in the previous section, are an example of a local basis, so splines can also straddle this category.
7.5.1 Fourier Spectral with Hacks
Figure 40: Preprocessing can make Fourier spectral filtering workable with aperiodic data. (a) Noisy data is padded with repeats of its end values. (b) The padded data is convolved with a moving average filter. (c) The original data is placed back in the center of the smoothed data from (b), and the combination is Fourier transformed. All modes except the lowest 50 are zeroed out, and the inverse transform is shown in red. Gibbs phenomenon is apparent but most extreme in the padded regions, which can be discarded. (d) The even extension of the data from (c) is Fourier transformed, filtered to the first 70 modes, and inverse transformed to produce the red curve. Gibbs phenomenon is much reduced.
Even though global Spectral Methods (section 4.2) are best applied in the absence of noise, due to Gibbs phenomenon in the case of the Fourier basis (section 4.2.1) and poor fits in the case of polynomial bases (section 4.2.2), noise is fundamentally such a frequency-dependent phenomenon (section 5), and the FFT is so attractively efficient, that fitting data with sinusoids is not uncommon. However, to get around some of the method’s weakness and work with arbitrary, likely-aperiodic data requires clever preprocessing.
The key is to recognize Gibbs phenomenon is caused by discontinuities and corners within a signal and at its edges (Figure 7), and is most profound near these triggers. One strategy, then, is to take the signal’s even extension (Figure 11), which gets rid of the discontinuity across a signal’s endpoints, but this can still result in a corner if the signal edges are not flat, and it comes at the cost of handling a double-length vector. An alternative approach, which can be used in tandem with even extension or independently, is to insulate the signal by padding its edges, being careful to smoothly transition between signal and padding so as not to subtly create more corners.
The preprocessed data (Figure 40) can then be Fourier transformed, ideal-low-pass filtered by zeroing out higher modes, optionally differentiated by multiplying by appropriate $ik$ (Equation 10), and inverse transformed. The smoothed signal or derivative is then found by slicing out part of the result.
Figure 41: Spectral derivatives of noisy cruise control data (Figure 22). At left, using no extension and no padding results in poor performance at the domain edges. Center, adding padding 100 samples long greatly improves the answer. Right, taking the even extension alone also works well.
7.5.2 Radial Basis Functions
Local functions with independent variable given by a radial distance, $r$, which evaluate to zero beyond some maximum distance, $\rho$—such as truncated Gaussians, $e^{-r^{2}/(2\sigma^{2})},\ r<\rho$—can also constitute a linearly independent basis, making the matrix in the linear inverse problem of Equation 5 banded, symmetric, and invertible in $O(N\rho^{2})$. We call this the $\mathbf{A}$ matrix, and visualize an example in the first panel of Figure 42. Unfortunately, $\mathbf{A}$ is typically ill-conditioned, having eigenvalues very close to 0, which makes coefficient solutions highly sensitive to noise in the target, $\mathbf{y}$. Unlike the Fourier case where sensitive higher modes, corresponding to tiny eigenvalues of $\mathbf{A}$, can be zeroed out, discarding any element of a local basis significantly deforms function reconstruction in some region, so we need to keep them all.
Figure 42: Visualizations of sparse, banded matrices used to fit radial basis functions to noisy data. Brighter colors indicate relatively higher values, and darker colors lower values. At left is a matrix filled with basis function samples, truncated when values get below 1e-4. This example has condition number about 1.2e9. Middle is a damped version of the same with smaller condition number, about 120. Right is a matrix filled with samples of the derivatives of basis functions.
One possible way to handle this dilemma is Tikhonov Regularization [Tikhonov], which attempts to keep coefficient values from blowing up by penalizing their $\ell_{2}$ norm:
$\underset{\mathbf{c}}{\min}\|\mathbf{A}\mathbf{c}-\mathbf{y}\|_{2}^{2}+\lambda\|\mathbf{c}\|_{2}^{2}$
where $\lambda$ is a chosen parameter. By taking the gradient w.r.t. $\mathbf{c}$ and setting equal to 0, the above is minimized by solving the linear inverse problem $(\mathbf{A}^{T}\mathbf{A}+\lambda\mathbb{I})\mathbf{c}=\mathbf{A}^{T}\mathbf{y}$. We can eigendecompose symmetric $\mathbf{A}=\mathbf{V}\boldsymbol{\Lambda}\mathbf{V}^{T}$ with $\mathbf{V}^{T}\mathbf{V}=\mathbb{I}$, and manipulate the solution into $(\boldsymbol{\Lambda}+\boldsymbol{\Lambda}^{-1}\lambda)\mathbf{c}^{\prime}=\mathbf{y}^{\prime}$, where $\mathbf{c}^{\prime}=\mathbf{V}^{T}\mathbf{c}$ and $\mathbf{y}^{\prime}=\mathbf{V}^{T}\mathbf{y}$ are just vector rotations. Because the matrices are diagonal, we can find $c_{i}^{\prime}=y_{i}^{\prime}/(\lambda_{i}+\lambda/\lambda_{i})$, where $\lambda_{i}$ is the $i^{\text{th}}$ eigenvalue from $\boldsymbol{\Lambda}$. In words, this means that for large eigenvalues, rotated coefficients are approximately proportional to rotated data divided by those eigenvalues, and as eigenvalues tend toward zero, coefficients do likewise.
Alternatively, to take full advantage of sparsity and avoid calculating $\mathbf{A}^{T}\mathbf{A}$, we can address ill conditioning by “damping” $\mathbf{A}$ itself, shifting all its eigenvalues up away from 0 by the addition of a scaled identity matrix:
$(\mathbf{A}+\mathbb{I}\lambda)\mathbf{c}=\mathbf{y}$
To see what this is doing relative to Tikhonov, we again eigendecompose to find $(\boldsymbol{\Lambda}+\mathbb{I}\lambda)\mathbf{c}^{\prime}=\mathbf{y}^{\prime}$, which has component-wise solution $c_{i}^{\prime}=y_{i}^{\prime}/(\lambda_{i}+\lambda)$. In words, large eigenvalues cause coefficients to be inversely proportional as before, but as $\lambda_{i}\to 0$, $c_{i}^{\prime}\to\frac{1}{\lambda}$ rather than 0 as for Tikhonov or $\infty$ in the undamped case.
This second scheme can also be thought of as manipulating the basis functions themselves, topping every “hill” with a “tower” to reach noisy data points. To obtain a smoothed estimate, we raze the towers and keep only the contributions of the original basis functions, using the classic reconstruction sum (Equation 3):
$\hat{x}(t)=\sum_{k=0}^{N-1}c_{k}\xi_{k}(t),\quad\xi_{k}(t)=\begin{cases}e^{-\frac{(t_{k}-t)^{2}}{2\sigma^{2}}}&|t_{k}-t|<\rho\\ 0&\text{otherwise}\end{cases}$
or in vector form: $\mathbf{\hat{x}}=\mathbf{A}\mathbf{c}$.
From this point, differentiation is extremely easy, because the differential operator is linear and can therefore be moved inside the sum: Just perform the reconstruction with $\dot{\xi}_{k}$, or in vector form, fill $\mathbf{\dot{A}}$ with samples from basis functions’ derivatives, as visualized in the third panel of Figure 42, and find $\mathbf{\hat{\dot{x}}}=\mathbf{\dot{A}}\mathbf{c}$.
Figure 43: Radial basis function fit derivatives of noisy cruise control data (Figure 22) with three different choices of hyperparameters.
7.6 Total Variation Regularization
Another way to think about noise in discrete series is that it causes extra disagreement between subsequent values, beyond what is necessary to trace the underlying true signal. We can measure the average distance between consecutive samples of a vector $\mathbf{v}$ with normalized Total Variation:
$\text{TV}(\mathbf{v})=\frac{1}{N}\|\mathbf{v}_{0:N-1}-\mathbf{v}_{1:N}\|_{1}$
One way to quell noise, then, is to drive down this quantity, an approach first introduced in [OsherFatemi92] in the context of noisy images and later applied by [Chartrand11] to the derivative estimation problem. Penalizing a noisy derivative’s total variation, in balance with its fidelity to the data, turns the ill-posed estimation problem into a well-posed one, i.e. regularizes it. We can describe this with the following optimization:
(44) $\underset{\mathbf{\hat{x}}}{\text{minimize}}\ \|\mathbf{y}-\mathbf{\hat{x}}\|_{2}^{2}+\gamma\text{TV}\Big(\frac{d^{\nu}\mathbf{\hat{x}}}{dt^{\nu}}\Big)$
where $\mathbf{y}$ are noisy data, $\mathbf{\hat{x}}$ is an estimate of the smoothed signal, $\gamma$ balances the terms, $\nu$ is the order of the derivative we wish to flatten, and the derivative, which cannot literally be taken on a vector of samples, is found as a vector of Finite Differences. With the insertion of Finite Difference to make the cost precise, the minimization problem becomes convex [^25] We could also maintain convexity while obtaining some outlier-robustness by modifying the first term of Equation 44 to be a Huber loss (Equation 30) instead of a sum of squares. and is directly solvable with tools like CVXPY [cvxpy].
Figure 44: Total Variation Regularized derivatives on noisy cruise control data (Figure 22). Notice the piecewise step, linear, and quadratic character as regularized derivative order increases.
Because the second term in Equation 44 is essentially an $\ell_{1}$ penalty, the solution is encouraged to have mostly zeros in its $\nu^{\text{th}}$ finite difference derivative, especially as $\gamma$ grows. This tends to result in a sparse selection of step changes in the $\nu^{\text{th}}$ derivative, giving TVR derivatives a particular character: piecewise constant for $\nu=1$, piecewise linear for $\nu=2$ (the integral of piecewise constant $2^{\text{nd}}$ derivative), and piecewise quadratic for $\nu=3$. Examples are shown in Figure 44. These stark shapes can optionally be convolved with a Gaussian kernel to soften corners.
7.7 Kalman Smoothing with a Naive Model
If a model is unknown, then it is possible to assume a model that, like Total Variation Regularization (TVR, section 7.6), attempts to stabilize the $\nu^{\text{th}}$ derivative, where $\nu$ is some order, and then apply model-based techniques (section 6) to estimate the derivative.
For example, if we assume a system model of the form:
$\mathbf{x}_{n}$ $=\mathbf{A}\mathbf{x}_{n-1}+\mathbf{B}\mathbf{u}_{n}+\mathbf{w}_{n},\ \ \mathbb{E}[\mathbf{w}\mathbf{w}^{T}]=\mathbf{Q}$
$\mathbf{y}_{n}$ $=\mathbf{C}\mathbf{x}_{n}+\mathbf{v}_{n},\ \ \mathbb{E}[\mathbf{v}\mathbf{v}^{T}]=\mathbf{R}$
then we can stabilize the second derivative by choosing:
$\mathbf{x}_{n}$ $=[x_{n,0},x_{n,1},x_{n,2}]^{T}=[\text{smooth sample},\ 1^{\text{st}}\text{ deriv. sample},\ 2^{\text{nd}}\text{ deriv. sample}]^{T}$
[…]
$\mathbf{A}=\begin{bmatrix}1&\Delta t&\frac{\Delta t^{2}}{2}\\ 0&1&\Delta t\\ 0&0&1\end{bmatrix},\quad\mathbf{B}=\begin{bmatrix}0\\ 0\\ 0\end{bmatrix},\quad\mathbf{C}=\begin{bmatrix}1&0&0\end{bmatrix}$
This performs Taylor integrations for the smoothed data and first derivative estimates, while holding the second derivative, sometimes called “acceleration” regardless of what the data represents, constant. Noisy observations flow into the compatible first dimension of the state and do not directly affect the latent derivatives.
Unlike in TVR, where stability of the $\nu^{\text{th}}$ derivative is maximized, the amount of stability in Kalman smoothing is squishy, depending on the noise level in the corresponding dimension of the process noise covariance matrix, $\mathbf{Q}$. If we assume process noise forces solely the last dimension, acting indirectly on its integrals over time, and note the measurement noise covariance matrix, $\mathbf{R}$, is $1\times 1$, then the covariance matrices can be treated as two scalar hyperparameters, $q$ and $r$:
$\mathbf{Q}=q\begin{bmatrix}\frac{\Delta t^{5}}{20}&\frac{\Delta t^{4}}{8}&\frac{\Delta t^{3}}{6}\\ \frac{\Delta t^{4}}{8}&\frac{\Delta t^{3}}{3}&\frac{\Delta t^{2}}{2}\\ \frac{\Delta t^{3}}{6}&\frac{\Delta t^{2}}{2}&\Delta t\end{bmatrix},\quad\mathbf{R}=\begin{bmatrix}r\\ \end{bmatrix}$
where the entries of the discrete-time $\mathbf{Q}$ matrix perform integration similar to $\mathbf{A}$. See section 8.3.
Figure 45: The frequency response of a constant-derivative RTS smoother is essentially yet another low-pass filter. Raising $q$ relative to $r$ shifts the peak rightward and broadens it out.
We can now solve the model-based Maximum A Priori (MAP) optimization problem from before (Equations 27 and 29) to produce a smoothed signal and derivatives. In fact, the $\ell_{2}$ minimizer is yet another explicit low-pass filter, with response [^26] To derive, choose $\Delta t$, $q$, and $r$; iterate 4 to convergence to find steady-state gain, $\mathbf{K}_{ss}$; form the closed-loop discrete LTI system, $G_{\text{cl}}=((\mathbb{I}-\mathbf{K}_{ss}\mathbf{C})\mathbf{A},\mathbf{K}_{ss},\mathbf{C},0)$ using standard $(\mathbf{A},\mathbf{B},\mathbf{C},\mathbf{D})$ notation [robust-control]; turn this into a transfer function with the formula $G(z)=\mathbf{C}(z\mathbb{I}-\mathbf{A})^{-1}\mathbf{B}+\mathbf{D}$ [robust-control], and find its frequency response. A similar process for the RTS backward pass (section 6.3) yields $\mathbf{L}_{ss}$, which, by treating the whole Kalman state as input and RTS state as output, can be used with Equation 28 to form the Multiple-Input Multiple-Output (MIMO) system $G_{\text{cl,RTS}}=(\mathbf{L}_{ss},\mathbb{I}-\mathbf{L}_{ss}A,\mathbb{I},0)$, whence we can extract a frequency response for the first dimension. The response of the two applied in sequence is the product. plotted in Figure 45.
Figure 46: Constant-derivative-model RTS smoothing derivatives of noisy cruise control data (Figure 22) with three different choices of hyperparameters. Derivatives are plucked directly from the second dimension of the state.
Just as in the model-based case (section 6.4), we can also optimize the MAP problem with alternative distance metrics like the $\ell_{1}$ norm or Huber loss to attain desired qualities like sparse solutions or robustness to outliers, as demonstrated in Figure 47, although this adds Huber-$M$ hyperparameters. This kind of flexibility is uncommon among smoothing methods; of those covered in this review, only spline and TVR optimization are similarly poised for loss function swap.
Figure 47: Constant-derivative-model RTS ($\ell_{2}$ norm loss) and Robust (Huber loss) solutions to the Kalman MAP problem (Equation 29) on noisy cruise control data with outliers (blue points in left panel). Derivatives are optimized for RMSE directly, demonstrating how an “accurate” fit to outliers can corrupt the smoothed estimate and derivative. Performance deteriorates less with the Huber loss, because it is better able to ignore outliers by less intensely penalizing their neglect. The solution is not as accurate as that in Figure 23, because the model does not match the true underlying system.
7.8 Performance Comparison
This section compares smoothing-based numerical differentiation methods, examining relative costs and advantages.
7.8.1 Computational Complexity
The first consideration is computational cost, addressed in Table 4. Each method has some complexity in itself, [^27] Computational complexities given in Table 4 assume equispaced data. Run times can be worse for irregularly-spaced data. See section 8. which usually takes the form $O(N\cdot f(\Phi))$, where $\Phi$ is the set of hyperparameters, interpreted in each method’s explanation above and in PyNumDiff code documentation [pynumdiff]. The only exception is Fourier Spectral, which is log-linear in $N$. Also of note is the size of each method’s hyperparameter set, $|\Phi|$, because although the choice of hyperparameters is in principle reduced to the single $\gamma$ from Equation 35 or 36, governing smoothness, we still need to run several Nelder-Mead optimizations to find the optimal $\Phi$, and this is more challenging in higher dimension.
Table 4: Computational complexity and hyperparameter set sizes of smoothing methods.
Method Complexity $\mathbf{|\Phi|}$
Filtering $\rightarrow$ Finite Difference $O(N\cdot\text{\footnotesize window size})$ 3
Iterated Finite Difference $O(N\cdot\text{\footnotesize order}\cdot\text{\footnotesize iterations})$ 2
Sliding Polyfit $O(N\frac{\text{(window size)}^{3}}{\text{stride}})$ 4
Savitzky-Golay $O(N\cdot\text{\footnotesize window size}\cdot\text{\footnotesize smoothing size})$ 3
Splines $O(N\cdot\text{\footnotesize degree}^{2}\cdot\text{\footnotesize iterations})$ 3
Fourier on Extension $O(N\log N)$ 3
Radial Basis Fit $O(N\rho^{2})$ 2
Total Variation Regularization $O(N\nu^{2}\log(\frac{1}{\epsilon}))$a 2
RTS Smoothing $O(N\nu^{3})$ 2b
Robust MAP Smoothing $O(N\nu^{3}\log(\frac{1}{\epsilon}))$c 5
  • a
    Equation 44 is “strongly convex” due to the quadratic term and can thus be practically solved with OSQP [boyd-osqp], which casts to a Quadratic Program and uses the ADMM algorithm to achieve “linear convergence” [michael-i-jordan-admm], meaning each iteration reduces error by a fixed multiplicative factor $\in\!(0,1)$, reaching error $\leq\!\epsilon$ in logarithmic steps [tibshirani2019convex]. Each step of OSQP entails a matrix inversion, but due to only local interactions of variables in the finite difference formulation (section 7.6), problem matrices are banded, with only $O(\nu)$ nonzero diagonals, so cost is $O(N\nu^{2})$.
  • b
    Scaling the two noise hyperparameters, $q$ and $r$ (section 7.7), by the same factor does not change the RTS frequency response (Figure 45), because the center of the joint distribution found at the Kalman filter’s final step (Figure 21) is invariant to absolute scale [pynumdiff]. Thus only the ratio between $q$ and $r$ makes a difference. Note this simplification is not possible for generalized MAP smoothing, because the absolute scales of $q$ and $r$ interact with distinct process and measurement Huber-$M$s.
  • c
    Piecewise linear quadratic (PLQ) losses like $\ell_{1}$ and Huber can be optimized by specialized interior point methods based on KKT conditions or by algorithms that approximate with smooth losses [aravkin, aravkin3]. When applied to measurements in series, the full system is sparse as in RTS (section 6.3), allowing linear per-iteration cost in $N$, rather than cubic cost as for general Second Order Cone Programs [socp-matmul, socp-complexity-mosek, tibshirani2019convex].
7.8.2 Accuracy and Bias
The other major consideration is method accuracy in terms of RMSE (Equation 33) and bias in terms of Error Correlation (Equation 34) against the true, withheld derivative. We are interested in whether the choice of method should change with different data characteristics: outliers, noise type, noise level, and the size of (consistent) $\Delta t$. We are also interested in whether the idiosyncrasies of particular data series favor particular differentiation algorithms, and we would like to know how sensitive methods are to the choice of cutoff frequency (bandlimit), which affects the smoothness of the optimized answers through the loss function’s $\gamma$, set according to the heuristic in Equation 37. In the presence of outliers, we optimize the robust loss function in Equation 36 with Huber parameter $M=2$, and in the absence of outliers, choose M=6M=6.282828We have observed empirically that, under conditions of Gaussian noise, optimizing the robust loss with M=6M\!=\!6 essentially always produces identical answers to optimizing the original RMSE-based loss, Equation 35. This makes sense theoretically, because inlier errors beyond $6\sigma$ from the mean error are exceedingly rare: 2*(1-scipy.stats.norm.cdf(6)) = 1.97e-9. Lowering $M$ to 2 provides a slight performance edge in the presence of outliers, allowing them to be weighted lighter, but in the absence of outliers M=2M\!=\!2 does begin to degrade performance—yet still only very slightly, because the portion of inliers within $2\sigma$ is more than 95%.
To examine these questions, we use six simulated examples, shown in Figure 48, each exhibiting its own unique attributes. Two are chosen by shape, to be composed of periodic or piecewise linear components, and the others arise from systems: two linear, as might be encountered in control theory, one nonlinear and chaotic, and one from mathematical biology. All simulations are on a similar scale so signal-to-noise ratio is comparable, and all can be readily recomputed with different step sizes.
Figure 48: Six simulations are used to produce noisy data and true underlying signal, as well as true derivatives (not shown). In the pictured examples, $t\in[0,4)$, $\Delta t=0.01$, and noise is Gaussian, $\eta\sim\mathcal{N}(\mu=0,\sigma=0.1)$. Both $\Delta t$ and $\eta$ are varied to test differentiation methods.
Figure 49: Noise can be sampled from a variety of probability density functions. For the purposes of our experiments, we use these three, as well as their stretched cousins.
We add noise from one of three different distributions, visualized in Figure 49: Normal Gaussian with $\mu,\sigma\!=\!(0,0.1)$, Laplace with $\mu,b\!=\!(0,0.1)$, or Uniform $\in[-0.2,0.2]$. The widths of these distributions are varied by multiplying their parameters by a noise scale. We further try corrupting a randomly-chosen 1% of samples with outlier values sampled uniformly from $\pm[50\%,150\%]$ of the max – min of their data series (e.g. left panel of Figure 47).
To determine a reasonable frequency cutoff, we examine signals’ power spectra in Figure 50 and observe a drop-off as in Figure 18. This occurs more rapidly for some than for others, but the same general shape holds, even with other choices of noise type, noise scale, and $\Delta t$. We choose to center our examinations of frequency cutoff around 3 Hz.
Figure 50: The power spectra of data $\mathbf{y}=\mathbf{x}+\boldsymbol{\eta},\ \boldsymbol{\eta}\sim\mathcal{N}(\mu=0,\sigma=0.1)$, calculated as the positive frequencies of $10\log_{10}|\text{FFT}(\mathbf{y})|^{2}$. Spectrum extends up to the Nyquist frequency $\frac{f_{s}}{2}=\frac{1}{2\Delta t}$, which for $\Delta t=0.01$ is 50 Hz.
The analysis space now has 7 dimensions: differentiation method, simulation, step size, noise type, noise scale, whether there are outliers, and cutoff frequency. There is also a bonus dimension: To get a sense of performance center and scatter, we need several rounds of simulation followed by optimization with different random seeds. Some dimensions must be explored exhaustively: 12 methods, 6 simulations, and we turn the number of random seeds all the way up to 52. To save compute and target our analysis at specific questions, we sweep along slices of the remaining orthotope, only varying one parameter at a time. We choose slices to all pass through a common central point, {cutoff frequency=3 Hz, $\Delta t$=0.01, noise type=normal, noise scale=1, outliers=False}), as symbolized in Figure 51.
$|\Delta t\text{s}|=4$$|\text{whether outliers}|=2$$|\text{noise scales}|=4$$|\text{frequency cutoffs}|=5$$|\text{noise types}|=3$
Figure 51: Notional look at slices in the performance analysis parameter space. Figures 5256 each correspond to one of the lines sketched here. The common central point is {cutoff frequency=3 Hz, $\Delta t$=0.01, noise type=normal, noise scale=1, outliers=False}.
Figures 5256 show RMSE and Error Correlation along each of these slices, with simulations marked by color, differentiation methods differentiated by symbol, and parameter choices separated into boxes. A method’s mean performance across the 52 randomly-seeded runs is marked by its symbol’s location, with bars indicating sample standard deviation. [^29] Confidence intervals on the mean can be calculated as a fraction of sample standard deviation, in this case about 0.278 for 95%, per scipy.stats.t.ppf(1-0.05/2, N-1)/np.sqrt(N) with $N=52$. For implementations of all methods, simulations, and the optimization loop used in this section, see PyNumDiff [pynumdiff]. Methods are plotted left to right in order of exposition in the text, but for clarity: KernelDiff refers to applying a smoothing kernel followed by FD, and SmoothAccelTVR refers to second-derivative-stabilizing TVR followed by convolution with a Gaussian kernel to dull sharp corners.
Figure 52: Outliers generally cause worse RMSE and Error Correlation.
Figure 53: Noise type does not have a profound effect on RMSE or Error Correlation, although the Laplace distribution, with the heaviest tails, causes slightly worse performance.
Figure 54: Increasing noise scale degrades average RMSE and Error Correlation simultaneously and increases variability for all methods. Scale=1 corresponds to $\mathcal{N}(0,0.1).$
Figure 55: Larger $\Delta t$ are difficult to handle for all methods.
Figure 56: Simulations without sharp corners continue to achieve lower RMSE as the cutoff frequency is decreased, favoring more smoothness during optimization. But this comes at the cost of sharply increasing Error Correlation, unless the underlying signal truly has very little frequency content.
7.9 Recommendations
The most unmistakable pattern in Figures 5256 is that all performances improve or degrade together with the difficulty of the problem, where difficulty is governed by:
  1. 1.
    the presence or absence of outliers,
  2. 2.
    the noise scale,
  3. 3.
    the size of $\Delta t$, because faster sampling raises the Nyquist frequency and allows better band separation of signal from noise, and
  4. 4.
    the smoothness of the true signal, because a spectrum that falls off more sharply, like that of the Logistic Growth simulation, allows more aggressive mitigation of high frequency noise without cutting into the signal.
In our experiments, optimizing robust loss, Equation 36 rather than Equation 35, greatly improves all methods’ performance in the presence of outliers. Only one method, RobustDiff, is specifically designed to handle outliers, by pairing an assumed model (section 7.7) with techniques from robust estimation (section 6.4). It achieves consistently lower bias than other methods and achieves the best accuracy for some but not all simulations (Figure 52). Unfortunately, with more hyperparameters than any other, RobustDiff is also the trickiest and most expensive to optimize. Demonstrating robustness of a different kind, PolyDiff (section 7.4.1) degrades less than other methods as the distance between samples is increased (Figure 55).
At lower bandwidth, noise scale, and step size, and even in the presence of outliers, TVRDiff (section 7.6) does noticeably best on the Triangles simulation, because its derivatives can match the piecewise character of the true derivative. But aside from this case, the best performer for a given simulation is typically inconsistent across sweeps and hardly distinguished from the pack. This is good news, because it means data series do not tend to favor particular methods for unknown reasons. Rather, sophisticated methods tend to do equally well, with only special cases favoring specific algorithms.
Remaining distinctions are less obvious. No method distinguishes itself as better able to handle a specific type of noise (Figure 53), and all methods are similarly challenged by increasing noise scale (Figure 54). Methods are likewise similarly sensitive to the choice of cutoff frequency (Figure 56). The choice of 3 Hz bandlimit is vindicated, because increasing $f$ (seeking less smooth derivatives) causes accuracy to worsen, while decreasing $f$ (seeking smoother derivatives) causes substantially worsening bias for all simulations except Logistic Growth, which can be explained by its very little frequency content (Figure 50).
To further compare methods, auxiliary considerations are summarized in Figure 57. These factors emphasize the value of algorithm flexibility, which can be underappreciated in theoretical discussion but is often important in practice. Indeed, one of the reasons neural networks have become so dominant is their endless capacity for recombination into architectures tailored to multivariate, multidimensional, sequential, incomplete, and structured data.303030“Time cannot wither nor custom stale her infinite variety.” –Shakespeare on the perceptron We recommend RTSDiff (section 7.7) as a general-purpose method for its outstanding versatility and narrowly superior accuracy in many experiments.
[…]
Figure 57: Ability of methods to handle auxiliary considerations. Black tildes mean neither yes nor no, often that a method could be extended to handle a situation, but at the cost of computational efficiency or mathematical or code simplicity. The heavy dependency at issue is a set of convex solvers, which, although easy to install and use in modern form [cvxpy], are based on underlying compilers and iterative algorithms that can fail to solve efficiently or at all, a source of potential brittleness.
8 Irregularly-Spaced Samples
We have made reference to a consistent step size, $\Delta x$ or $\Delta t$, throughout this review, but in the context of real data, sample spacing is anything but guaranteed and can even be highly variable. This is a wrinkle worth addressing, lest differentiation be practically limited to a fraction of experimental data streams.
Because each differentiation method is built from distinct mathematics, some handle irregular spacing more naturally than others. Especially in cases where performance may not be all that decisive (Figure 57), this can have massive bearing on method choice. Table 5 summarizes which methods discussed in this review can be made to handle irregular steps and why, and subsequent subsections take a closer look at how several methods natively handle or can be extended to handle this situation.
Table 5: How amenable differentiation methods are to variable step size, $\widetilde{\Delta t}$.
Method $\mathbf{\widetilde{\Delta t}}$? Notes
AutoDiff N/A Not for differentiating data samples, rather for sampling fixed derivatives. But can be sampled wherever.
[…]
Fourier Spectral $\checkmark$ There is the Non-Uniform FFT for this [nufft]. With even extension/padding, can work for noisy, aperiodic data.
Chebyshev X No speed advantage if not cosine-spaced samples, and bad basis for separating noise.
Finite Elements N/A Perhaps confusingly, works with irregular $\Delta t$ internally, but uses governing equation as input rather than samples, except optionally for boundary conditions.
Kalman Methods $\checkmark$ Evaluating $e^{M},\ M\in\mathbb{R}^{m\times m}$ costs $O(m^{3})$ [matrix-exp], so same asymptotic complexity, though more involved math.
Smoothing $\rightarrow$ FD X Filters do not behave consistently with irregular $\Delta t$.
Sliding Polyfit $\checkmark$ Use sliding window in index space rather than dependent-variable space to capture same-sized groups of points across the domain.
Savitzky-Golay X Ceases to be a consistent filter, so essentially back to sliding polyfit.
Splines $\checkmark$ Can be solved as efficiently as uniform case, and placing knots is no more complicated than before.
Radial Basis Fit $\checkmark$ Local basis functions can be plopped over data anywhere, although they interact less if further separated.
TVR X Complicated FD equations turn the cost function gnarly.
8.1 Finite Difference with Irregular Steps
There is nothing stopping us from solving the stencil Vandermonde inverse problem of Equation 2 with irregular stencils. We do this by multiplying both sides by $\Delta x$ to put stencil locations back in units of the independent variable and then using whichever fractional or decimal stencil locations, $s_{0},...,s_{S-1}]$, match data locations.
For example, using a “centered” stencil on three points with distances $d_{01}$ between the first two and $d_{12}$ between the last two gives:
$y^{\prime}_{n}\approx$ $\ c_{0}y_{n-1}+c_{1}y_{n}+c_{2}y_{n+1}$
$=$ $\ c_{0}[y_{n}-d_{01}\cdot y^{\prime}_{n}+\frac{(-d_{01})^{2}}{2!}y^{\prime\prime}_{n}+O(d_{01}^{3})]+c_{1}y_{n}$
$+c_{2}[y_{n}+d_{12}\cdot y^{\prime}_{n}+\frac{d_{12}^{2}}{2!}y^{\prime\prime}_{n}+O(d_{12}^{3})]$
$\rightarrow$ $\begin{bmatrix}1&1&1\\ -d_{01}&0&d_{12}\\ \frac{d_{01}^{2}}{2}&0&\frac{d_{12}^{2}}{2}\end{bmatrix}\begin{bmatrix}c_{0}\\ c_{1}\\ c_{2}\end{bmatrix}=\begin{bmatrix}0\\ 1\\ 0\end{bmatrix}$
Solving bespoke linear inverse problems like this incurs an extra $O(\text{order}^{3})$ cost per data point. We should also be mindful that error bounds degrade, because $\|\mathbf{c}\|_{2}\leq 1/\sigma_{\text{min}}$, and the smallest singular value of the matrix, $\sigma_{\text{min}}$, is maximized with equal distances. In addition, if iterating this process (section 7.3), we have to be careful to use correct trapezoid widths when integrating, and handle stencils adaptively at the endpoints to avoid coefficients with magnitude over 1, which amplify noise.
8.2 Splines with Irregular Steps
Continuous B-splines (Figure 38) and their analytic derivatives can be sampled anywhere in the domain, and knots can be added anywhere, so splines naturally handle irregularly-spaced data. Wherever samples are taken, there are always only $O(d)$ nonzero basis functions at each, where $d$ is the degree of the spline, so the matrix $\mathbf{B}$ from Equation 43 is always sparse, $\mathbf{B}^{T}\mathbf{B}$ remains banded, and the fit problem is just as easy to solve.
8.3 Kalman Smoothing with Irregular Steps
Our exposition of Kalman filtering in section 6.1 uses a discrete-time model of system dynamics, which is convenient when thinking in terms of computation, but we can imagine this model is actually arising from the continuous evolution of a system over constant increments of time[axelssonandgustafsson, crassidis_junkins]. We can write the underlying continuous model as:
$\frac{d\mathbf{x}}{dt}=\mathcal{A}\mathbf{x}(t)+\mathcal{B}\mathbf{u}(t)+\mathbf{w}(t)$
where script $\mathcal{A}$ and $\mathcal{B}$ are the continuous-time state evolution and control matrices, respectively, and $\mathbf{x}$, $\mathbf{u}$, and $\mathbf{w}$ now vary continuously. Note that sampled observations are still discrete and obey $\mathbf{y}(t_{n})=\mathbf{C}\mathbf{x}(t_{n})+\mathbf{v}_{n},\ \mathbf{v}_{n}\sim\mathcal{N}(0,\mathbf{R})$, just as before.
We would like to know $\mathbf{x}(t_{n}+\Delta t)$ for some time in the future, so we solve the above with the method of integrating factor. We make the simplifying assumption that $\mathbf{u}$ is constant over this interval, so-called “zero-order hold”, which is valid if control inputs or synchronized data streams only get updates at sampling times, a reasonable assumption if they come from a computed control or sensing loop. We substitute $\tau=t-t_{n}$ because all terms will become integrands, and integrating from 0 simplifies the math.
$\frac{d\mathbf{x}}{d\tau}=\mathcal{A}\mathbf{x}(\tau)+\mathcal{B}\mathbf{u}+\mathbf{w}(\tau)\rightarrow\underbrace{e^{-\mathcal{A}\tau}\frac{d\mathbf{x}}{d\tau}-e^{-\mathcal{A}\tau}\mathcal{A}\mathbf{x}(\tau)}_{=\frac{d}{d\tau}\Big(e^{-\mathcal{A}\tau}\mathbf{x}(\tau)\Big)}=e^{-\mathcal{A}\tau}\big(\mathcal{B}\mathbf{u}+\mathbf{w}(\tau)\big)$
$\quad\quad\quad\rightarrow\underbrace{\int\limits_{0}^{\Delta t}\frac{d}{d\tau}\Big(e^{-\mathcal{A}\tau}\mathbf{x}(\tau)\Big)d\tau}_{\mathclap{e^{-\mathcal{A}\tau}\mathbf{x}(\tau)\Big|_{0}^{\Delta t}=e^{-\mathcal{A}\Delta t}\mathbf{x}(\Delta t)-\mathbb{I}\mathbf{x}(0)}}=\int\limits_{0}^{\Delta t}e^{-\mathcal{A}\tau}\big(\mathcal{B}\mathbf{u}+\mathbf{w}(\tau)\big)d\tau$
$\quad\quad\quad\rightarrow e^{-\mathcal{A}\Delta t}\mathbf{x}(\Delta t)=\mathbf{x}(0)+\int\limits_{0}^{\Delta t}e^{-\mathcal{A}\tau}\mathcal{B}d\tau\ \mathbf{u}+\int\limits_{0}^{\Delta t}e^{-\mathcal{A}\tau}\mathbf{w}(\tau)d\tau$
$\quad\quad\quad\rightarrow\mathbf{x}(\Delta t)=e^{\mathcal{A}\Delta t}\mathbf{x}(0)+\int\limits_{0}^{\Delta t}e^{\mathcal{A}(\Delta t-\tau)}\mathcal{B}d\tau\ \mathbf{u}+\underbrace{\int\limits_{0}^{\Delta t}e^{\mathcal{A}(\Delta t-\tau)}\mathbf{w}(\tau)d\tau}_{\mathbf{w}_{d}(\Delta t)\sim\mathcal{N}}$
where $\mathbf{w}_{d}$ is discrete noise. The middle term can be simplified a bit with a change of variables: Letting $\tilde{\tau}=\Delta t-\tau$, then when $\tau=0,\ \tilde{\tau}=\Delta t$, and when $\tau=\Delta t,\ \tilde{\tau}=0$, and $\frac{d\tilde{\tau}}{d\tau}=-1\rightarrow d\tau=-d\tilde{\tau}$.
$\int\limits_{0}^{\Delta t}e^{\mathcal{A}(\Delta t-\tau)}\mathcal{B}d\tau=\int\limits_{\Delta t}^{0}e^{\mathcal{A}\tilde{\tau}}\mathcal{B}(-d\tilde{\tau})=\int\limits_{0}^{\Delta t}e^{\mathcal{A}\tilde{\tau}}\mathcal{B}d\tilde{\tau}=\mathbf{B}(\Delta t)$
where the bold symbol denotes the discrete-time matrix.
The noise term, $\mathbf{w}_{d}$, is a zero-mean Gaussian random variable because integrating $\mathbf{w}(\tau)$ against $e^{\mathcal{A}(\Delta t-\tau)}$, which is just another matrix, constitutes a linear functional, which preserves the mean and Gaussian character of the process, only stretching its covariance. We can find this new covariance as:
(45) $\mathbb{E}[\mathbf{w}_{d}(\Delta t)\mathbf{w}_{d}(\Delta t)^{T}]-\overset{\text{\normalsize 0}}{\cancel{\mathbb{E}[\mathbf{w}_{d}]\mathbb{E}[\mathbf{w}_{d}^{T}]}}=\mathbb{E}\Big[\int\limits_{0}^{\Delta t}e^{\mathcal{A}(\Delta t-\tau)}\mathbf{w}(\tau)\mathbf{w}(\tau)^{T}e^{\mathcal{A}^{T}(\Delta t-\tau)}d\tau\Big]$
$\quad\quad\quad=\int\limits_{0}^{\Delta t}e^{\mathcal{A}(\Delta t-\tau)}\underbrace{\mathbb{E}[\mathbf{w}(\tau)\mathbf{w}(\tau)^{T}]}_{\mathclap{\text{constant, continuous-time }\mathcal{Q}}}e^{\mathcal{A}^{T}(\Delta t-\tau)}d\tau=\int\limits_{0}^{\Delta t}e^{\mathcal{A}\tilde{\tau}}\mathcal{Q}e^{\mathcal{A}^{T}\tilde{\tau}}d\tilde{\tau}=\mathbf{Q}(\Delta t)$
where $\mathcal{Q}$ is constant because it is the expected value of a stationary random matrix, which is deterministic, irrespective of $\tau$, and in the last step we have done the same $\tilde{\tau}=\Delta t-\tau$ trick as before.
We can now find the estimated state and error covariance for time offset $\Delta t$ by taking expected values:
(46) $\mathbf{\hat{x}}(\Delta t)$ $=\mathbb{E}[\mathbf{x}(\Delta t)]=\mathbb{E}[e^{\mathcal{A}\Delta t}\mathbf{x}(0)+\mathbf{B}(\Delta t)\mathbf{u}+\mathbf{w}_{d}(\Delta t)]$
$=e^{\mathcal{A}\Delta t}\mathbb{E}[\mathbf{x}(0)]+\mathbb{E}[\mathbf{B}(\Delta t)\mathbf{u}]+\mathbb{E}[\mathbf{w}_{d}(\Delta t)]=e^{\mathcal{A}\Delta t}\mathbf{\hat{x}}(0)+\mathbf{B}(\Delta t)\mathbf{u}$
(47) $\mathbf{P}(\Delta t)$ $=\mathbb{E}\big[\big(\mathbf{x}(\Delta t)-\mathbb{E}[\mathbf{x}(\Delta t)]\big)\big(\mathbf{x}(\Delta t)-\mathbb{E}[\mathbf{x}(\Delta t)]\big)^{T}\big]$
$=\mathbb{E}\big[\big(e^{\mathcal{A}\Delta t}\mathbf{x}(0)+\cancel{\mathbf{B}(\Delta t)\mathbf{u}}+\mathbf{w}_{d}(\Delta t)-e^{\mathcal{A}\Delta t}\mathbb{E}[\mathbf{x}(0)]-\cancel{\mathbf{B}(\Delta t)\mathbf{u}}-\mathbb{E}[\mathbf{w}_{d}(\Delta t)]\big)\big(\text{ditto}\big)^{T}\big]$
$=\mathbb{E}\big[\big(e^{\mathcal{A}\Delta t}\big(\mathbf{x}(0)-\mathbb{E}[\mathbf{x}(0)]\big)+\big(\mathbf{w}_{d}(\Delta t)-\mathbb{E}[\mathbf{w}_{d}(\Delta t)]\big)\big)\big(\text{ditto}\big)^{T}\big]$
$=e^{\mathcal{A}\Delta t}\underbrace{\mathbb{E}\big[\big(\mathbf{x}(0)-\mathbb{E}[\mathbf{x}(0)]\big)\big(\mathbf{x}(0)-\mathbb{E}[\mathbf{x}(0)]\big)^{T}\big]}_{P(0)}e^{\mathcal{A}^{T}\Delta t}+\overset{\text{\normalsize 0}}{\cancel{\text{cross terms}}}+\mathbf{Q}(\Delta t)$
where cross terms are 0 because noise is assumed to be uncorrelated with state estimation error.
By treating any $\mathbf{x}_{n},\mathbf{P}_{n}$ as initial conditions through the time shift $\tau=t-t_{n}$, we can use Equations 46 and 47 to find $\mathbf{x}_{n+1},\mathbf{P}_{n+1}$ for arbitrary step size. All we have left to do is compute discrete-time $\mathbf{A}(\Delta t)=e^{\mathcal{A}\Delta t}$, $\mathbf{B}(\Delta t)$, and $\mathbf{Q}(\Delta t)$ given continuous-time $\mathcal{A}$, $\mathcal{B}$, and $\mathcal{Q}$. The state evolution matrix exponential can be taken directly, but the discrete control matrix and process noise covariance involve integrals. Thankfully, there is a very elegant trick [integration-via-expM, axelssonandgustafsson, crassidis_junkins] to integrate by exponentiating a special matrix, which allows us to solve for all these quantities at once.
First consider the generic block triangular matrix:
$\mathcal{M}=\begin{bmatrix}\mathcal{M}_{1}&\mathcal{M}_{2}\\ 0&\mathcal{M}_{3}\end{bmatrix}$
Taking a matrix exponential is equivalent to plugging the matrix in to the Taylor expansion:
$e^{\mathcal{M}t}=\mathbb{I}+\mathcal{M}t+\frac{\mathcal{M}^{2}t^{2}}{2!}+\frac{\mathcal{M}^{3}t^{3}}{3!}+\cdots$
The sum of integer powers of a block triangular matrix remains block triangular, so the result has structure:
$\mathbf{M}(t)=\begin{bmatrix}\mathbf{M}_{1}(t)&\mathbf{M}_{2}(t)\\ 0&\mathbf{M}_{3}(t)\end{bmatrix}$
We can thus set up a differential equation:
$\frac{d}{dt}e^{\mathcal{M}t}=\mathcal{M}e^{\mathcal{M}t}=\mathcal{M}\mathbf{M}(t)$
$\rightarrow\frac{d}{dt}\begin{bmatrix}M_{1}&M_{2}\\ 0&M_{3}\end{bmatrix}=\begin{bmatrix}\mathbf{\dot{M}}_{1}&\mathbf{\dot{M}}_{2}\\ 0&\mathbf{\dot{M}}_{3}\end{bmatrix}=\begin{bmatrix}\mathcal{M}_{1}&\mathcal{M}_{2}\\ 0&\mathcal{M}_{3}\end{bmatrix}\begin{bmatrix}\mathbf{M}_{1}(t)&\mathbf{M}_{2}(t)\\ 0&\mathbf{M}_{3}(t)\end{bmatrix}$
where over-dots represent the derivative, which can be safely distributed to each block independently. By equating blocks, we can turn this into three equations:
$\mathbf{\dot{M}}_{1}=\mathcal{M}_{1}\mathbf{M}_{1},\quad\mathbf{\dot{M}}_{2}=\mathcal{M}_{1}\mathbf{M}_{2}+\mathcal{M}_{3}M_{3},\quad\mathbf{\dot{M}}_{3}=\mathcal{M}_{3}\mathbf{M}_{3}$
These can be solved by integrating factor. Blocks 1 and 3 have the same governing equation, so we show the solution procedure only of block 1:
$\mathbf{\dot{M}}_{1}=\mathcal{M}_{1}\mathbf{M}_{1}\rightarrow\mathbf{\dot{M}}_{1}-\mathcal{M}_{1}\mathbf{M}_{1}=0\rightarrow e^{-\mathcal{M}_{1}\tau}\mathbf{\dot{M}}_{1}-e^{-\mathcal{M}_{1}\tau}\mathcal{M}_{1}\mathbf{M}_{1}=0$
$\rightarrow\frac{d}{d\tau}\big(e^{-\mathcal{M}_{1}\tau}\mathbf{M}_{1}(\tau)\big)=0\rightarrow\int\limits_{0}^{t}\frac{d}{d\tau}\big(e^{-\mathcal{M}_{1}\tau}\mathbf{M}_{1}\big)d\tau=\int\limits_{0}^{t}0d\tau$
$\rightarrow e^{-\mathcal{M}_{1}\tau}\mathbf{M}_{1}(\tau)\Big|_{0}^{t}=0\rightarrow e^{-\mathcal{M}_{1}t}\mathbf{M}_{1}(t)-\mathbb{I}\underbrace{\mathbf{M}_{1}(0)}_{\mathclap{\text{on-diagonal block of }e^{\mathcal{M}0}=\mathbb{I}}}=0\rightarrow\mathbf{M}_{1}(t)=e^{\mathcal{M}_{1}t}$
For block 2:
$\mathbf{\dot{M}}_{2}-\mathcal{M}_{1}\mathbf{M}_{2}=\mathcal{M}_{2}\mathbf{M}_{3}\rightarrow\int\limits_{0}^{t}\frac{d}{d\tau}(e^{-\mathcal{M}_{1}\tau}\mathbf{M}_{2})d\tau=\int\limits_{0}^{t}e^{-\mathcal{M}_{1}\tau}\mathcal{M}_{2}\mathbf{M}_{3}d\tau$
$\rightarrow e^{-\mathcal{M}_{1}\tau}\mathbf{M}_{2}(\tau)\Big|_{0}^{t}=e^{-\mathcal{M}_{1}t}\mathbf{M}_{2}(t)-\mathbb{I}\underbrace{\mathbf{M}_{2}(0)}_{\mathclap{0\text{ because off-diagonal in }e^{\mathcal{M}0}}}$
$\rightarrow\mathbf{M}_{2}(t)=\int\limits_{0}^{t}e^{\mathcal{M}_{1}(t-\tau)}\mathcal{M}_{2}\mathbf{M}_{3}(\tau)d\tau=\int\limits_{0}^{t}e^{\mathcal{M}_{1}\tilde{\tau}}\mathcal{M}_{2}\mathbf{M}_{3}(t-\tilde{\tau})d\tilde{\tau}$
where in the last step we use the same variable transformation as before. Substituting the solution for $\mathbf{M}_{3}(t)$, we get:
$\mathbf{M}_{2}(t)=\int\limits_{0}^{t}e^{\mathcal{M}_{1}\tilde{\tau}}\mathcal{M}_{2}e^{\mathcal{M}_{3}(t-\tilde{\tau})}d\tilde{\tau}$
Notice that the solution for blocks 1 and 3 looks like the basic matrix exponentiation we have to do to find $\mathbf{A}$, and the solution for block 2 looks suspiciously like the integral expression for $\mathbf{Q}$ from Equation 45. Let’s choose $\mathcal{M}_{1}=\mathcal{A}$, $\mathcal{M}_{2}=\mathcal{Q}$, and $\mathcal{M}_{3}=-\mathcal{A}^{T}$. We then form the matrix:
$\mathcal{M}=\begin{bmatrix}\mathcal{A}&\mathcal{Q}\\ 0&-\mathcal{A}^{T}\end{bmatrix}$
and exponentiate to produce:
$\mathbf{M}(\Delta t)=\begin{bmatrix}e^{\mathcal{A}\Delta t}&\int\limits_{0}^{\Delta t}e^{\mathcal{A}\tilde{\tau}}\mathcal{Q}e^{\mathcal{A}^{T}(\tilde{\tau}-\Delta t)}d\tilde{\tau}\\ 0&e^{-\mathcal{A}^{T}\Delta t}\end{bmatrix}=\begin{bmatrix}\mathbf{A}(\Delta t)&\mathbf{Q}(\Delta t)\mathbf{A}\!^{-T}(\Delta t)\\ 0&\mathbf{A}\!^{-T}(\Delta t)\end{bmatrix}$
where -T denotes the inverse transpose. Similarly, choosing $\mathcal{M}_{1}=\mathcal{A}$, $\mathcal{M}_{2}=\mathcal{B}$, and $\mathcal{M}_{3}=0$ yields:
$\mathbf{M}(\Delta t)=\begin{bmatrix}\mathbf{A}(\Delta t)&\mathbf{B}(\Delta t)\\ 0&\mathbb{I}\end{bmatrix}$
So we can simply pick out a couple of the matrices we need, and find the other by multiplication since $\mathbf{A}\!^{-T}\mathbf{A}^{T}=e^{-\mathcal{A}^{T}\Delta t}e^{\mathcal{A}^{T}\Delta t}=\mathbb{I}$.
For all the math that got us here, this is very short to code and fairly cheap to run: For a model of dimension $\nu$ the matrix $\mathcal{M}\in\mathbb{R}^{2\nu\times 2\nu}$ costs $O(\nu^{3})$ to exponentiate [matrix-exp], but both the Kalman filter (4) and RTS smoother (Equation 28) are already $O(\nu^{3})$ per iteration due to dense matrix multiplications or inversions, so variable time steps are slightly more complicated and expensive but do not actually worsen the asymptotic complexity.
RTS smoothing extends to irregular steps just as easily as the Kalman filter, because it simply takes the history of all a priori and a posteriori $\mathbf{x}$ and $\mathbf{P}$ estimates and combines distributions, agnostic to whether they come from Equations 46 and 47 or from the update equations in 4. The exponentiation causes covariances $\mathbf{Q}$ and $\mathbf{P}$ to grow continuously between steps, but the optimal rule for combining Gaussians remains the same regardless of their size.
9 Conclusions
Numerical differentiation is a fundamental operation that appears across all scientific and engineering domains, yet in practice selecting an appropriate method requires careful consideration of problem structure, data characteristics, and computational constraints. This review has organized the landscape of differentiation algorithms into five major scenarios: static, analytic relationships best handled by Automatic Differentiation; regular and irregular simulations where Spectral Methods and Finite Elements excel, respectively; and cases of noisy data that bifurcate into model-based approaches leveraging Kalman filtering frameworks versus model-free smoothing techniques. The key insight is that more constrictive assumptions enable better performance—periodic signals unlock the power of Fourier spectral methods; known system dynamics enable optimal Kalman smoothing; and even naive constant-derivative models provide remarkably effective regularization.
For practitioners facing noisy data without prior models, our comprehensive experimental comparison reveals that while sophisticated methods achieve similar accuracy under ideal conditions, auxiliary considerations often prove decisive. RTSDiff (section 7.7) emerges as a versatile general-purpose choice, offering narrow performance advantages alongside exceptional flexibility for variable step sizes and data interpolation. For specific challenges, PolyDiff (section 7.4) handles large time steps more gracefully, while RobustDiff provides superior outlier rejection at the cost of substantially increased optimization computation. Particular narratives may naturally suit alternative choices, such as ButterDiff where frequency plays a central role or TVRDiff (section 7.6) to identify piecewise derivatives. One can even apply multiple methods to the same data to reveal challenging sections where methods disagree or to corroborate hyperparameter choices.
Different end uses of derivative estimates can dictate which methods, and which hyperparameters to choose. For example, when derivatives are used for control, a smoother derivative, at the cost of higher error correlation and RMSE, may be a better choice to prevent high control efforts [crassidis_junkins]. Meanwhile, for data-driven modeling applications, low error correlations may be preferred at the cost of higher RMSE to prevent the overemphasis of low frequency dynamics [sindy]. In most cases, however, a balanced estimate will likely be a good starting point. The loss function framework presented in Equations 3536, combined with the cutoff frequency heuristic of Equation 37, enables principled hyperparameter optimization that consistently approaches Pareto-optimal solutions across methods and problem types, reducing the practitioner’s burden from navigating dozens of method-specific parameters to selecting a single smoothness parameter informed by signal bandwidth.
Ultimately, successful numerical differentiation demands matching method assumptions to problem characteristics. As data-driven science continues to expand, robust and flexible derivative estimation remains part of the essential computational infrastructure, and understanding the tradeoffs between accuracy, computational cost, and applicability enables researchers to extract maximum insight from their measurements and simulations.
Acknowledgments
We are especially grateful to Sasha Aravkin for extensive discussions on convex optimization techniques related to robust and nonsmooth formulations. We would also like to thank Jing Yu for lending her clarity and depth to guide us through some of the complex relationships among robust control methods. And finally we recognize Alexey Yermakov for thoroughly proofreading an early draft. The authors acknowledge support from the National Science Foundation AI Institute in Dynamic Systems (grant number 2112085).